### v1

# Cospectral vertices, Walk-regular planar graphs and the echolocation problem

Abstract.
We study cospectral vertices on finite graphs in relation to the echolocation problem on Riemannian manifolds. First, We prove a computationally simple criterion to determine whether two vertices are cospectral. Then, we use this criterion in conjunction with a computer search to find minimal examples of various types of graphs on which cospectral but non-similar vertices exist, including minimal walk-regular non-vertex-transitive graphs, which turn out to be non-planar. Moreover, as our main result, we classify all finite walk-regular planar graphs, proving that such graphs must be vertex-transitive.
1. Introduction
Let $G$ be a finite graph with adjacency matrix $A$. Let $\phi_{1},\phi_{2},\ldots,\phi_{n}$ be an orthonormal basis of eigenfunctions (eigenvectors) of $A$ with respective eigenvalues $\lambda_{1},\lambda_{2},\ldots,\lambda_{n}$. We say two vertices $a$ and $b$ are cospectral if
(1.1) $\sum_{\lambda_{j}=\lambda}|\phi_{j}(a)|^{2}=\sum_{\lambda_{j}=\lambda}|\phi_{j}(b)|^{2}\qquad\text{ for all eigenvalues $\lambda$.}$
All graphs considered in this article will be finite, simple, undirected, and connected unless otherwise specified.
The study of cospectral vertices has gained more attention recently, (see e.g., [God12, GS17, ${}^{+}$19, EKL19, KSSW20]), especially due to their significant role in quantum information theory. It has been shown [God12] that for a perfect quantum state transfer to occur between two vertices in an interacting qubits network, it is necessary for these vertices to be cospectral.
We say two vertices $a$, $b$ are similar if there is an automorphism of $G$ mapping $a$ to $b$. Two similar vertices are clearly also cospectral. Recently, Kempton, Sinkovic, Smith, and Webb [KSSW20] show that two vertices are cospectral if and only if they are similar in a certain weaker sense. We are interested in the exceptional situation where two vertices are cospectral but not similar. We come to this problem from the smooth setting, where a graph is replaced with a Riemannian manifold. One can formulate a notion of cospectral points that reads exactly as (1.1), provided the adjacency matrix is replaced with the Laplace-Beltrami operator (see Remark 3.2). The question of whether or not there exists a pair of cospectral points that are not in the same orbit has a musical interpretation: Can one hear, up to symmetry, where a drum is struck? We refer to this question as the echolocation problem. The last two authors show that for most compact Riemannian manifolds, the answer is ‘yes’ [WX23]. So far, there are no known examples of connected manifolds where the answer is ‘no’.
In search of such negative examples, we turn to the graph-theoretic setting, where we might use a computer search. In fact, it is well-known that graphs with non-similar cospectral vertices exist. The first such example recorded in literature is a tree with 9 vertices observed by Schwenk [Sch73]. To find more such examples and study their properties, we must find a way to check when two vertices are cospectral that avoids floating point errors. The criterion (1.1) is not convenient in this way, so we need an alternative.
Our method builds off of work by Godsil and Smith [GS17], who show that two vertices $a$ and $b$ of a graph are cospectral with respect to the adjacent matrix $A$, if and only if $(A^{k})_{a,a}=(A^{k})_{b,b}$ for all $k\in\mathbb{N}.$ In other words, $a$ and $b$ are cospectral if, for all $k\in{\mathbb{N}}$, the number of walks of length $k$ starting and ending at $a$ is equal to the number of walks of length $k$ starting and ending at $b$. See also [CB23, Mon22] for results in other settings. When it comes to computation, this condition can be used to distinguish non-cospectral vertices but would require seemingly infinitely many operations to confirm that two vertices are cospectral. In this paper, we show that, in fact, one only needs to compute a finite number of powers of $A$ to check cospectrality for a finite graph.
Proposition 1.1.
Let $G$ be a graph (possibly with self-loops) with $n$ vertices and $A$ be the corresponding adjacent matrix. Then two vertices $a$ and $b$ are cospectral if and only if
(1.2) $(A^{k})_{a,a}=(A^{k})_{b,b},\quad\text{for all }0\leq k\leq n-1.$
Moreover, this condition is sharp in the sense that the upper limit $n-1$ cannot be made smaller.
We remark that, even though Godsil and Smith did not state it explicitly, the first part of Proposition 1.1 follows by the proof of [GS17, Lemma 2.1]. We shall include a proof of Proposition 1.1 in Section 2 for completeness. Then, we employ this condition in a computer search to find minimal examples of regular graphs with cospectral yet non-similar vertices in Section 3.
Theorem 1.2.
The smallest graph containing a pair of vertices that are cospectral but not similar has eight vertices.
Riemannian manifolds have some rigid structure that graphs, in general, lack. We impose a variety of kinds of regularity on our graphs to emulate this structure. We first search amongst the regular graphs for examples. This setting comes with the benefit that the adjacency matrix $A$ and the two standard symmetric normalizations of the Laplacian $L=D-A$ and $\mathcal{L}=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}$ all have precisely the same eigenfunctions. Indeed, we find such examples, but only amongst graphs of ten or more vertices.
Theorem 1.3.
The smallest regular graph containing a pair of non-similar cospectral vertices has ten vertices.
Even more structured are walk-regular graphs [GM80], all of whose vertices are cospectral. A vertex $a$ has degree $(A^{2})_{a,a}$, and hence all walk-regular graphs are also regular. Seeking minimal examples amongst walk-regular graphs yields:
Theorem 1.4.
The smallest walk-regular graph containing a pair of non-similar vertices has twelve vertices.
Smooth analogs of walk-regular graphs were studied by Wang and the last two authors in [WWX23]. There we show that every compact two-dimensional surface satisfying the smooth version of walk-regularity is also homogeneous—the isometry group acts transitively on the points in the surface. The assumption that our manifold is two-dimensional is essential to the proof, and it is yet unknown if there are any higher-dimensional manifolds that are walk-regular but not homogeneous. In addition, none of the walk-regular graphs we found are planar. This leads us to consider planar walk-regular graphs.
We are able to prove the following discrete analog of the result of [WWX23].
Theorem 1.5.
Every walk-regular planar graph with degree $d\geq 3$ is isomorphic to the polyhedral graphs of one of the following:
  • 5 regular polyhedrons, also known as Platonic solids: tetrahedron, hexahedron, octahedron, dodecahedron, and icosahedron;
  • 13 semi-regular polyhedrons, also known as Archimedean solids, pictured in Figure 1;
  • the $m$-gonal prisms for $m=3$ and $m\geq 5$; and
  • the $m$-gonal antiprisms for $m\geq 4$.
Consequently, every walk-regular planar graph is vertex-transitive.
A direct consequence of the classical theorem of Mani [Man71] says that all planar vertex-transitive finite graphs with degree at least 3 are exactly those listed above. Herbert and Wilfried [FI79] employed modified Schläfli symbol to give a direct proof of this fact. Theorem 1.5 is a refinement of this result of Mani and Herbert–Wilfried, as walk-regularity is, in general, strictly weaker than vertex-transitivity. The modified Schläfli symbol again plays a crucial role in our proof of Theorem 1.5.
Figure 1. All the 13 semi-regular polyhedrons. The graphics are due to Wikipedia user Cyp [Cyp12].
Organization of the paper
In Section 2, we prove Proposition 1.1. In Section 3, we give a short overview of our computer search and present a few of our minimal examples. The code in full is available on GitHub along with a repository of minimal examples (https://github.com/ewy-man/graph-echolocation). Section 4 is dedicated to the proof of Theorem 1.5.
Acknowledgements
Xi was supported by the National Key Research and Development Program of China No. 2022YFA1007200 and NSF China Grant No. 12171424. Kong was supported by NSF China Grant No. 12301109 and the Fundamental Research Funds for the Central Universities of China No. YJ202343. Wyman was supported by NSF grant DMS-2204397. The second listed author extends a thanks to Richard Lange for his very patient guidance through the process of setting up a GitHub repository.
2. Closed walks of bounded length determine if two vertices are cospectral
Let $A$ be the adjacency matrix of some graph $G$ on $n$ vertices, or any real-symmetric $n\times n$ matrix for that matter. Let $\phi_{1},\ldots,\phi_{n}$ be a real orthonormal eigenbasis for $A$ with respective eigenvalues $\lambda_{1},\ldots,\lambda_{n}$. For each vertex $a$, consider the measure
$\mu_{a}=\sum_{j}|\phi_{j}(a)|^{2}\delta_{\lambda_{j}}$
on ${\mathbb{R}}$. Observe that two vertices $a$ and $b$ are cospectral if and only if $\mu_{a}=\mu_{b}$.
Let $m\leq n$ denote the count of distinct eigenvalues and note each measure $\mu_{a}$ is determined by its pairing with degree $m-1$ polynomials. In particular, the measure $\mu_{a}$ is uniquely determined by its pairings
[…]
with the standard basis $\{1,t,\ldots,t^{m-1}\}$ for the degree $m-1$ polynomials. We can arrive at the same set of quantities from a different direction, namely
$A_{a,a}^{k}=\sum_{j,\ell}\phi_{j}(a)(\phi_{j}^{t}A^{k}\phi_{\ell})\phi_{\ell}(a)=\sum_{j}|\phi_{j}(a)|^{2}\lambda_{j}^{k}.$
In summary, we have shown:
Lemma 2.1.
Let $A$ be a real-symmetric $n\times n$ matrix. Then, $\mu_{a}=\mu_{b}$ if and only if
$A^{k}_{a,a}=A^{k}_{b,b}\qquad\text{ for all $k=0,1,\ldots,n-1$.}$
We note that the number $n-1$ of counts is sharp, since the adjacency matrix $A$ of a simple path with $n$ vertices for every odd $n=2\ell+1$ (i.e., $A_{i,j}=\delta_{|i-j|=1}$) as well as the adjacency matrix $A$ of a simple path with $n$ vertices and a self-loop attached to the last vertex for every even $n=2\ell$ (i.e., $A_{i,j}=\delta_{|i-j|=1}+\delta_{i=j=n}$) satisfies
$A^{k}_{\ell,\ell}=A^{k}_{\ell+1,\ell+1}\quad\text{for all}~{}k=0,1,\ldots,n-2\quad\text{but}\quad A^{n-1}_{\ell,\ell}<A^{n-1}_{\ell+1,\ell+1}.$
3. Minimal examples
An exhaustive list of minimal examples of graphs with cospectral, non-similar vertices under regularity conditions of varying strength can be found on GitHub (https://github.com/ewy-man/graph-echolocation) along with the code used to generate them.
The code is written in the Julia language (see [BEKS17, ${}^{+}$21]) and contains a number of tools:
  1. (1)
    A few functions that generate exhaustive lists of non-isomorphic graphs on a specified number of vertices. An optional filter may be passed to these functions to narrow down and speed up the search.
  2. (2)
    A function which returns the orbits of the action of the isomorphism group on the vertices.
  3. (3)
    A function that determines the cospectrality of the vertices via Proposition 1.1.
  4. (4)
    A number of other convenience functions, such as those that determine if a graph is walk-regular, or if there exist non-similar cospectral pairs.
Once loaded into the Julia REPL, these functions can be called to generate the minimal examples below and to check that they are indeed minimal. The first listed operation is the most time-intensive, but typically does not take more than a minute to run on a personal computer.
3.1. Minimal graphs with non-similar cospectral vertices
The smallest graphs that contain non-similar cospectral pairs of vertices have eight vertices. Of all 12346 non-isomorphic graphs on 8 vertices, 126 of them contain non-similar cospectral vertices. Four examples are depicted in Figure 2. Furthermore, all of the 126 minimal examples are connected and none of them are trees.
Figure 2. Four examples of graphs with eight vertices with a pair of non-similar cospectral vertices. The non-similar cospectral pairs are marked in red.
3.2. Minimal regular graphs with non-similar cospectral vertices
The smallest regular graphs which contain non-similar cospectral pairs of vertices have ten vertices, and of those only those of degree 3, 4, 5, and 6 have minimal examples.
$d$ number of $d$-regular examples number of $d$-regular graphs
$3$ 3 21
$4$ 22 60
$5$ 22 60
$6$ 3 21
The symmetry in the table is not a coincidence.
Proposition 3.1.
Let $G=(V,E)$ be a finite graph and let $G^{\prime}=(V,E^{c})$ denote its edge complement.
  1. (1)
    $G$ and $G^{\prime}$ have the same graph isomorphism group.
  2. (2)
    If $G$ and $G^{\prime}$ are both connected, then two vertices are cospectral in $G$ if and only if they are cospectral in $G^{\prime}$.
The proposition explains the symmetry we see in the table. Indeed, all of the minimal examples counted by the table here happen to be connected. Each degree $3$ example’s edge complement is a minimal degree $6$ example, and similarly for degree $4$ and $5$ examples.
Remark 3.2.
Given a $d$-regular graph $G$ with adjacency matrix $A$, its graph Laplacian is written
$L=dI-A.$
As such, $A$ and $L$ share precisely the same eigenfunctions. Hence, two vertices $a$ and $b$ of a $d$-regular graph are cospectral if and only if
$\sum_{\lambda_{j}=\lambda}|\phi_{j}(a)|^{2}=\sum_{\lambda_{j}=\lambda}|\phi_{j}(b)|^{2}$
where $\phi_{1},\ldots,\phi_{n}$ is an orthonormal Laplace eigenbasis with $L\phi_{j}=\lambda_{j}\phi_{j}$.
Proof.
Part (1) is obvious and well-known. For part (2), note the Laplacian $L^{\prime}$ on $G^{\prime}$ can be written in terms of the Laplacian $L$ on $G$ as
$L^{\prime}=nI-\mathbf{1}-L,$
where $\mathbf{1}$ is the $n\times n$ matrix with a one in every entry. It suffices to show that an eigenfunction $\phi$ for $L$ is also an eigenfunction for $L^{\prime}$. Since $G$ and $G^{\prime}$ are both connected, their trivial eigenspaces are both spanned by the constant function, so it suffices to show $L$ and $L^{\prime}$ share nonconstant eigenfunctions. If $\phi$ is an eigenfunction of $L$ with eigenvalue $\lambda>0$, it is orthogonal to the constant eigenspace and hence lies in the kernel of $\mathbf{1}$. Hence,
$L^{\prime}\phi=(n-\lambda)\phi.$
The reverse argument is identical. ∎
We remark that, as a consequence, the edge complement of a walk-regular graph is again walk-regular. This is Theorem 4.2 in the work of Godsil and Mckay [GM80] on walk-regular graphs.
3.3. Minimal walk-regular, non-vertex-transitive graphs
There are no walk-regular, non-vertex-transitive graphs which have less than 12 vertices. There are exactly four minimal examples with twelve vertices, one each of degree 4, 5, 6, and 7. Again, these minimal examples are connected, and so can be put into edge-complement pairs; the examples of degrees $4$ and $7$ are edge complements, and so are the examples of $5$ and $6$. The degree $4$ and $5$ examples are depicted in Figure 3.
Refer to caption      Refer to caption
Figure 3. These graphs and their edge-complements are the only walk-regular, non-vertex-transitive graphs on $12$ vertices. Red vertices belong to one orbit and blue vertices belong to the other.
4. Walk-regular planar graphs
Here we will set up the tools we will need in order to classify the planar walk-regular graphs and prove Theorem 1.5.
[…]
$r(v)=(r_{1}(v),r_{2}(v),\ldots,r_{d}(v)),$
where $r_{i}(v)$ is the number of edges of $f_{i}(v)$. The first face $f_{1}(v)$, as well as the labeling direction, are chosen to minimize the label $r(v)$ lexicographically so that it is uniquely defined. In particular, $3\leq r_{1}(v)=\min_{1\leq i\leq d}r_{i}(v)$. With such a label, we count the total number $f$ of faces in the graph by
$f=\sum_{v}p(v),\qquad\text{where}\ \ p(v):=\sum_{i=1}^{d}\frac{1}{r_{i}(v)}.$
Using Euler’s formula $n+f-e=2$ and $e=nd/2$, we get
(4.1) $\frac{1}{n}\sum_{v}p(v)=\frac{d-2}{2}+\frac{2}{n}$
We let $k(v)=\#\{1\leq i\leq d:r_{i}(v)=r_{1}(v)\}$. An easy observation is
Lemma 4.1.
In a walk-regular planar graph with $d\geq 3$,
[…]
Proof.
First we let $r_{1}=\min_{v}r_{1}(v)$. Then it follows from (4.1) that
$\frac{d}{r_{1}}\geq\frac{d-2}{2}+\frac{2}{n}.$
Thus $r_{1}\in\{3,4,5\}$ for $d=3$, and $r_{1}=3$ for $d=4$ or $5$. Let $v_{1}$ be a vertex such that $r_{1}(v_{1})=r_{1}$. Suppose that there is a vertex $v_{2}$ that satisfies $r_{1}(v_{2})>r_{1}$. If $r_{1}\in\{3,5\}$, we have
$A^{r_{1}}_{v_{2},v_{2}}=0<2k(v_{1})=A^{r_{1}}_{v_{1},v_{1}}.$
If $r_{1}=4$, then $d=3$ and
$A^{4}_{v_{2},v_{2}}=15<15+2k(v_{1})=A^{4}_{v_{1},v_{1}}.$
In any case, $v_{1}$ and $v_{2}$ are not cospectral. Hence $r_{1}(v)\equiv r_{1}$ provided that the graph is walk-regular. By a similar closed walk counting argument, we conclude that
$A_{v,v}^{r_{1}}-A_{v^{\prime},v^{\prime}}^{r_{1}}=2k(v)-2k(v^{\prime})$
for any pair of vertices $v,v^{\prime}$, and thus $k(v)\equiv k$ by walk-regularity. ∎
In order to prove Theorem 1.5, we shall prove the following lemma.
Lemma 4.2.
For any walk-regular planar graph with degree $d\geq 3$, the modified Schläfli symbol is constant over all vertices, that is
(4.2) $r(v)\equiv(r_{1},r_{2},\cdots,r_{d})\qquad\textit{for all vertices}~{}v.$
However, the proof of this lemma appears to be quite delicate. The 3-regular planar graph, as illustrated in Figure 5, features cospectral vertices $v_{1}$ and $v_{2}$ with $r(v_{1})=(3,5,5)$ but $r(v_{2})=(3,5,6)$. Despite vertex $v_{1}$ having two neighboring pentagons and $v_{2}$ only one, the adjacent triangle $\{v_{5},v_{7},v_{9}\}$ contributes two additional closed walks of length 5 to $v_{2}$. Moreover, this graph is ‘almost’ walk-regular, with all vertices except $v_{4}$ being cospectral. Therefore, establishing (4.2) succinctly appears challenging. Instead, we justify (4.2) in each individual case.
We remark that, once we prove Lemma 4.2, it is routine to check that Theorem 1.5 holds in each case. Indeed, the numbers of vertices, edges, and faces are determined by Lemma 4.2 and Euler’s identity. The graph can then be reconstructed in a deterministic fashion, starting with an arbitrary polygon as the outer polygon. We illustrate this fact for the most complicated case when $r(v)\equiv(3,4,5,4)$ in Figure 4. The only exception is the case when $r(v)\equiv(3,4,4,4)$, for which one runs into two possible reconstructions, the polyhedral graphs of the $(3,4,4,4)$-solid and twisted $(3,4,4,4)$-solid. However, the twisted one is not walk-regular. See Case (13) in the proof.
Figure 4. Step by step reconstruction of the $(3,4,5,4)$-solid.
Figure 5. A 3-regular planar graph on 10 vertices.
Let $v_{1}$ be a vertex such that $p(v_{1})=\max_{v}p(v)$. In view of Lemma 4.1, to prove Lemma 4.2, we only need to discuss the sixteen cases enumerated below. For the convenience of the reader, we have also included an exhaustive list of the possible walk-regular planar graphs belonging to each case. From now on and throughout the paper, for the purpose of simplicity, we shall not distinguish between a solid and its polyhedral graph.
Case $d,r_{1},k$ Solids
(1) 3, 3, 3 Tetrahedron
(2) 3, 3, 2
(3) 3, 3, 1 Triangular prism, $(3,6,6)$-, $(3,8,8)$-, and $(3,10,10)$-solids
(4) 3, 4, 3 Hexahedron
(5) 3, 4, 2 $m$-gonal prism, $m\geq 5$
(6) 3, 4, 1 $(4,6,6)$-, $(4,6,8)$-, and $(4,6,10)$-solids
(7) 3, 5, 3 Dodecahedron
(8) 3, 5, 2
(9) 3, 5, 1 $(5,6,6)$-solid
(10) 4, 3, 4 Octahedron
(11) 4, 3, 3 $m$-gonal antiprism, $m\geq 4$
(12) 4, 3, 2 $(3,4,3,4)$- and $(3,5,3,5)$-solids
(13) 4, 3, 1 $(3,4,4,4)$- and $(3,4,5,4)$-solids
(14) 5, 3, 5 Icosahedron
(15) 5, 3, 4 $(3,3,3,3,4)$- and $(3,3,3,3,5)$-solids.
(16) 5, 3, $\leq$ 3
Of these cases, (2) and (8) are simply impossible to realize. It is also impossible to realize case (16), since
$p(v)\leq\frac{1}{3}+\frac{1}{3}+\frac{1}{3}+\frac{1}{4}+\frac{1}{4}=\frac{3}{2}$
violates (4.1). One quickly sees the symbols $r(v)$ of the graphs realizing cases (1), (4), (7), (10), and (14) must be constant.
Next, we focus on cases (5), (11), (9), and (15), which are easier to resolve. The remaining cases (3), (6), (12), and (13) are more involved. We will treat these cases after.
Case (5)
($d=3,r_{1}=4,k=2$) For each $m=r_{3}(v_{1})\geq 5$, the face $f_{3}(v_{1})$ surrounded by $m$ quadrangular faces determines the unique pattern up to isomorphism. In this case, $r(v)\equiv(4,4,m)$ and the graph is isomorphic to the $m$-gonal prism.
Case (11)
($d=4,r_{1}=3,k=3$) For each $m=r_{4}(v_{1})\geq 4$, the face $f_{4}(v_{1})$ surrounded by $m$ triangular faces determines the unique pattern up to isomorphism. In this case, $r(v)\equiv(3,3,3,m)$ and the graph is isomorphic to the $m$-gonal antiprism.
Starting from the next case, we often need to count the closed walks of length $m$ for some $m\geq 4$ and two adjacent vertices $v,v^{\prime}$ with $r(v)=r(v_{1})\neq r(v^{\prime})$. Our strategy of counting is to consider only the $\lfloor\frac{m}{2}\rfloor$-neighborhood of $\{v,v^{\prime}\}$, which turns out to be determined by the values of $d,r_{1},k$ and $r(v_{1})$, with at most two possibilities.
Case (9)
($d=3,r_{1}=5,k=1$) Each face with $m\geq 6$ edges must be surrounded by exact $m/2$ pentagonal faces and thus $m$ is even. As $p(v_{1})>\frac{1}{2}$ by (4.1), the only possibility is $r(v_{1})=(5,6,6)$. Suppose $r(v)\not\equiv(5,6,6)$. Then there exists a pentagonal face with two adjacent vertices $v,v^{\prime}$ such that $r(v)=(5,6,6)$ and $r(v^{\prime})=(5,6,m)$ with $m>6$. The $3$-neighborhood of $\{v,v^{\prime}\}$ is unique, with 23 vertices (Figure 6). This gives
$A_{v,v}^{6}=91>89=A^{6}_{v^{\prime},v^{\prime}}$
Hence $r(v)\equiv(5,6,6)$, the graph is isomorphic to the $(5,6,6)$-solid.
Figure 6. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(5,6,6)$ and $r(v^{\prime})=(5,6,m)$ with $m>6$.
Case (15)
($d=5,r_{1}=3,k=4$) As $p(v_{1})>\frac{3}{2}$ by (4.1), $r_{5}(v_{1})=4$ or $5$. When $r_{5}(v_{1})=4$, we suppose $r_{5}(v)\not\equiv 4$. Then there exist two adjacent vertices $v,v^{\prime}$ such that $r_{5}(v)=4<r_{5}(v^{\prime})$. Counting closed walks of length $4$ in the $2$-neighborhood of $\{v,v^{\prime}\}$ (Figure 7), we have
$A_{v,v}^{4}=59>57=A_{v^{\prime},v^{\prime}}^{4}.$
Hence $r(v)\equiv(3,3,3,3,4)$, the graph is isomorphic to the $(3,3,3,3,4)$-solid.
Figure 7. Two possible patterns of the $2$-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,3,3,3,4)$ and $r(v^{\prime})=(3,3,3,3,m)$ with $m>4$.
In the case that $r_{5}(v_{1})=5$, $r_{5}(v)\not\equiv 5$, there must exist two adjacent vertices $v,v^{\prime}$ with $r_{5}(v)=5<r_{5}(v^{\prime})$. Counting closed walks of length 5 in the 2-neighborhood of $\{v,v^{\prime}\}$ (Figure 8) gives
$A_{v,v}^{5}=182>180=A_{v^{\prime},v^{\prime}}^{5}.$
Hence $r(v)\equiv(3,3,3,3,5)$, the graph is isomorphic to the $(3,3,3,3,5)$-solid.
Figure 8. Two possible patterns of the $2$-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,3,3,3,5)$ and $r(v^{\prime})=(3,3,3,3,m)$ with $m>5$.
We now move on to the more involved cases.
Case (3)
($d=3,r_{1}=3,k=1$) Denote the vertex set of the triangular face $f_{1}(v_{1})$ by $\{v_{1},v_{2},v_{3}\}$, where $v_{2}$ does not belong to the face $f_{2}(v_{1})$. Note that every face with $m\geq 4$ edges should be surrounded by exact $m/2$ triangular faces, and thus $m$ must be even. Suppose $r_{2}(v_{1})>10$. Then
$p(v_{1})\leq\frac{1}{3}+\frac{1}{12}+\frac{1}{12}=\frac{1}{2},$
which violates (4.1). Hence $r_{2}(v_{1})=4,6,8$ or $10$.
Sub-case: $r_{2}(v_{1})=4$
If $r_{3}(v_{1})>4$, it follows that $r_{2}(v_{2})>4$ and
$A_{v_{1},v_{1}}^{4}=17>15=A_{v_{2},v_{2}}^{4}.$
Hence the walk-regularity implies $r_{3}(v_{1})=4$. In this case, the only possibility is $r(v)\equiv(3,4,4)$, and the graph is isomorphic to the triangular prism.
Sub-case: $r_{2}(v_{1})=6$
Suppose $r_{3}(v_{1})>6$. Then $r_{2}(v_{2})>6$, otherwise we have $r(v_{3})=(3,6,6)$ that violates the minimality of $p(v_{1})$. Counting the closed walks with length $6$ in the $3$-neighborhood of $\{v_{1},v_{2}\}$ (Figure 9)
$A_{v_{1},v_{1}}^{6}=91>89=A_{v_{2},v_{2}}^{6}.$
Thus $r(v_{1})=(3,6,6)$ by the walk-regularity.
Figure 9. 3-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,6,m)$ with $m>6$.
Now we claim that $r(v)\equiv(3,6,6)$. Indeed, whenever this is false, there exists a triangular face with vertex set $\{v,v^{\prime},v^{\prime\prime}\}$ such that $r(v)=(3,6,6)$ and $r(v^{\prime})=(3,6,m)$ with $m>6$ (Figure 10), which gives
$A_{v,v}^{6}=93>91=A_{v^{\prime},v^{\prime}}^{6}.$
Hence $r(v)\equiv(3,6,6)$, and the graph is isomorphic to the $(3,6,6)$-solid.
Figure 10. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,6,6)$ and $r(v^{\prime})=(3,6,m)$ with $m>6$.
Sub-case: $r_{2}(v_{1})=8$
As in the previous sub-case, $r_{3}(v_{1})>8$ implies $r_{2}(v_{2})>8$, and a counting in the $4$-neighborhood of the pair $\{v_{1},v_{2}\}$ (Figure 11) provides
$A_{v_{1},v_{1}}^{8}=593>591=A_{v_{2},v_{2}}^{8},$
thus $r(v_{1})=(3,8,8)$.
Figure 11. 4-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,8,m)$ with $m>8$.
If $r(v)\equiv(3,8,8)$ is violated, there is a triangular face with vertices $v,v^{\prime}$ such that $r(v)=(3,8,8)$ and $r(v^{\prime})=(3,8,m)$ with $m>8$ (Figure 12), and
$A_{v,v}^{8}=595>593=A_{v^{\prime},v^{\prime}}^{8}.$
This shows that $r(v)\equiv(3,8,8)$, and the graph is isomorphic to the $(3,8,8)$-solid.
Figure 12. 4-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,8,8)$ and $r(v^{\prime})=(3,8,m)$ with $m>8$.
Sub-case: $r_{2}(v_{1})=10$
Again, $r_{3}(v_{1})>10$ implies $r_{2}(v_{2})>10$, and a counting in the 5-neighborhood of $\{v_{1},v_{2}\}$ (Figure 13) gives
$A_{v_{1},v_{1}}^{10}=4225>4223=A_{v_{2},v_{2}}^{10},$
thus $r(v_{1})=(3,10,10)$.
Figure 13. 5-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,10,m)$ with $m>10$.
If $r(v)\equiv(3,10,10)$ is violated, there is a triangular face with vertices $v,v^{\prime}$ such that $r(v)=(3,10,10)$ and $r(v^{\prime})=(3,10,m)$ with $m>10$ (Figure 14), and
$A_{v,v}^{10}=4227>4225=A_{v^{\prime},v^{\prime}}^{10}.$
This shows that $r(v)\equiv(3,10,10)$, and the graph is isomorphic to the $(3,10,10)$-solid.
Figure 14. 5-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,10,10)$ and $r(v^{\prime})=(3,10,m)$ with $m>10$.
Case (6)
($d=3,r_{1}=4,k=1$) Each face with $m\geq 5$ should be surrounded by exact $m/2$ quadrangular faces and thus $m$ is even. As $p(v_{1})>\frac{1}{2}$ by (4.1), we must have $r_{2}(v_{1})=6$ and $r_{3}(v_{1})=6,8$ or $10$.
Sub-case: $r_{3}(v_{1})=6$
Suppose $r(v)\not\equiv(4,6,6)$. Then there exists a quadrangular face with two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,6)$ and $r(v^{\prime})=(4,6,m)$ with $m>6$ (Figure 15). By counting closed walks we obtain
$A_{v,v}^{6}=115>113=A_{v^{\prime},v^{\prime}}^{6}.$
Thus $r(v)\equiv(4,6,6)$, and the graph is isomorphic to the $(4,6,6)$-solid.
Figure 15. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,6)$ and $r(v^{\prime})=(4,6,m)$ with $m>6$.
Sub-case: $r_{3}(v_{1})=8$
For this we claim that for every vertex $v$ with $r(v)=(4,6,8)$, its adjacent vertex $v^{\prime}$ outside the hexagonal face $f_{2}(v)$ also has $r(v^{\prime})=(4,6,8)$. Indeed, if this is violated by some adjacent pair $v,v^{\prime}$ (Figure 16), then
$A_{v,v}^{6}=113>111=A_{v^{\prime},v^{\prime}}^{6}.$
This verifies the claim.
Figure 16. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,8)$ and $r_{2}(v^{\prime})>6$.
Now suppose $r(v)\not\equiv(4,6,8)$. Then by the claim there exist two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,8)$ and $r(v^{\prime})=(4,6,m)$ with $m>8$ (Figure 17). A counting of closed walks with length $8$ gives
$A_{v,v}^{8}=811>809=A_{v^{\prime},v^{\prime}}^{8}.$
Hence $r(v)\equiv(4,6,8)$, and the graph is isomorphic to the $(4,6,8)$-solid.
Figure 17. 4-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,8)$ and $r(v^{\prime})=(4,6,m)$ with $m>8$.
Sub-case: $r_{3}(v_{1})=10$
Note that any two adjacent vertices $v,v^{\prime}$ with $r(v)=(4,6,10)$ and $r_{2}(v^{\prime})>6$ (Figure 18) satisfy
$A_{v,v}^{6}=113>111=A_{v^{\prime},v^{\prime}}^{6}.$
This shows that any vertex that is adjacent to some vertex with label $(4,6,10)$ should lie on a hexagonal face.
Figure 18. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,10)$ and $r_{2}(v^{\prime})>6$.
Moreover, if $r(v)\not\equiv(4,6,10)$, then there exist two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,10)$ and $r(v^{\prime})=(4,6,m)$ with $m>10$ (Figure 19), which give
$A_{v,v}^{10}=6065>6063=A_{v^{\prime},v^{\prime}}^{10}.$
Hence $r(v)\equiv(4,6,10)$, and the graph is isomorphic to the $(4,6,10)$-solid.
Figure 19. 5-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,10)$ and $r(v^{\prime})=(4,6,m)$ with $m>10$.
Case (12)
($d=4,r_{1}=3,k=2$) We start with a lemma.
Lemma 4.3.
Let $G$ be $d$-regular with $d\geq 3$. Then,
$A_{v_{i},v_{i}}^{4}-(2d^{2}-d)=2(\#\text{ of $4$-cycles through $v_{i}$}).$
Proof.
Every closed walk of length $4$ is either contractible or traces a $4$-cycle in one of two directions, and there are exactly $d(d-1)+d^{2}$ contractible closed walks of length $4$. ∎
We understand each vertex is adjacent to two triangles and two other faces, say of degree $4\leq m\leq m^{\prime}$ respectively, both larger than $3$. We subdivide the argument into four sub-cases where there are zero, one, two, or more than two $4$-cycles through each vertex.
Sub-case: Zero 4-cycles
There are no squares, and no two triangles share an edge. Such graphs have plenty of structure. Firstly, each face of $\geq 5$ sides is adjacent only to triangular faces. Furthermore, each such face is, through each vertex, opposed to another face of $\geq 5$ sides. We first establish that, in this situation,
$A_{v_{i},v_{i}}^{5}-60=2\times(\#\text{ of pentagons incident to $v_{i}$})$
Our job is now to eliminate the cases where
$r(v)=(3,5,3,6)\text{ or }(3,5,3,7)$
for each $v$, and we will be left with $r(v)\equiv(3,5,3,5)$. In fact, if we have either case above for one vertex, we have it for all of them. We can construct a $3$-neighborhood of the vertex $v$ (Figure 20) with either a hexagon or a face of $\geq 7$ sides and obtain
$A_{v,v}^{6}=270+\begin{cases}2&r(v)=(3,5,3,6)\\ 0&r(v)=(3,5,3,7).\end{cases}$
Figure 20. Two possible 3-neighborhood of $v$.
If we attempt to construct either graph with either $r\equiv(3,5,3,6)$ or $r\equiv(3,5,3,7)$, we quickly run into absurdities. Hence, we are left with the $(3,5,3,5)$ solid.
Sub-case: one $4$-cycle
First, we rule out the possibility that such a graph has a square face. If we start with a square face, we obtain a graph resembling that in Figure 21, which we will call a “bee” since it looks like it has a head with two antennae, a thorax with two wings, and an abdomen.
Figure 21. Our “bee” graph.
Note, one or both of the antennae may attach to a wing to form a pentagon. Running some software, we find that, for any vertex $v$ of the square,
$A_{v,v}^{5}=\begin{cases}68&\text{ if no pentagon is incident to $v$}\\ 70&\text{ if the remaining face incident to $v$ is a pentagon}.\end{cases}$
On the other hand, the vertex $v^{\prime}$ at the “neck” of the bee satisfies
$A_{v^{\prime},v^{\prime}}^{5}=62+2\times(\#\text{ of pentagons incident to $v^{\prime}$})\leq 66,$
and so we have $A_{v^{\prime},v^{\prime}}^{5}<A_{v,v}^{5}$, a contradiction.
We have shown that the only $4$-cycles we find are the boundary of the union of two triangular faces. Since there are no squares, there must be at least one pentagon. The graph self-destructs upon any attempt to draw a $1$-neighborhood of the pentagon.
Sub-case: two $4$-cycles
There are a few situations we must rule out. We first establish that each vertex must be incident to a square face. Suppose $v$ is not adjacent to a square face. Then, $r(v)=(3,3,m,m^{\prime})$ or $(3,m,3,m^{\prime})$, where $m,m^{\prime}\geq 5$. The first is incident to only one $4$-cycle, so is ruled out. We try to draw the graph starting at this vertex and obtain the one in Figure 22, which forces a vertex to have at least three $4$-cycles.
Figure 22. The left-center vertex is forced to be incident to three $4$-cycles.
With this in hand, any construction of a graph with a vertex with the label $(3,3,4,m)$ for $m\geq 5$ fails after drawing a few faces. Hence, every vertex of our graph must have label $(3,4,3,4)$, and we obtain the $(3,4,3,4)$-solid.
Sub-case: More than two $4$-cycles
The only way a vertex can be incident to four $4$-cycles is if its $2$-neighborhood is that in Figure 23. However, this vertex’s neighbors are only ever incident to three $4$-cycles.
Figure 23. $2$-neighborhood of a vertex incident to four $4$-cycles.
Hence, all that remains is to exclude the case where the vertices are incident to exactly three $4$-cycles.
A similar argument as the one above excludes the case where $v$ is incident to a face of $5$ or more sides. Hence, for each vertex $v$, we have either $r(v)=(3,4,3,4)$ or $r(v)=(3,3,4,4)$. There is a unique graph with these properties, depicted in Figure 24.
Figure 24. The graph with exactly three $4$-cycles through each vertex.
This graph, however, is not walk-regular. The three innermost and outermost vertices all have $74$ returning walks of length $5$, while the other vertices have $76$.
Case (13)
($d=4,r_{1}=3,k=1$) As in the proof of the previous case, we subdivide the argument according to the number of 4-cycles through each vertex. Since there is exactly 1 triangle face adjacent to each vertex, the number of 4-cycles is determined by the number of adjacent squares. Therefore, each vertex must be adjacent to the same number of squares. By (4.1), we must have $p(v_{1})>1$, and thus there must be at least two squares adjacent to each vertex.
Sub-case: three squares at each vertex
This case is straightforward since we already know that $r(v)\equiv(3,4,4,4)$. We remark that, in this case, when one wants to reconstruct the $(3,4,4,4)$-solid, there are two possible graphs, the other one being the twisted $(3,4,4,4)$-solid graph. However, there are vertices on the twisted $(3,4,4,4)$ solid that have different numbers of closed walks of length 7, and thus, this twisted case should be ruled out. See Figure 25.
Figure 25. (3,4,4,4)-solid and twisted (3,4,4,4)-solid.
Sub-case: two squares at each vertex
In this case, $r(v_{1})$ is either $(3,4,4,5)$ or $(3,4,5,4)$. We first consider the case $r(v_{1})=(3,4,4,5)$. One can try to draw such a graph starting from $v_{1}$ and end up with a neighborhood of the shape as in Figure 26.
Figure 26. $r(v_{1})=(3,4,4,5)$
Here, vertex $v_{1}$ is what we started with. We claim that the unknown polygon adjacent to vertex $v_{3}$ must be a pentagon. Then a similar argument will show that all outer polygons must be pentagons. To do so, we count the closed walks of length 5 for $v_{1}$ and for $v_{3}$. Non-trivial ones are either a pentagon or a triangle attached to a square, which we shall call a “house”. $v_{1}$ and $v_{3}$ always share the same number of “houses”, and we are done with the claim.
Next, we claim that we will get a contradiction if the center polygon is not a pentagon. In this case, for $v_{1}$ and $v_{2}$ to have the same number of closed walks of length 5, we cannot create more “houses” in the outer layer for vertex $v_{1}$. This forces us to get 3 squares in a row. After drawing all the squares we arrive at Figure 27, which is impossible since there are vertices adjacent to more than 2 triangles.
Figure 27. An impossible graph when the center is not a pentagon.
Therefore, we conclude that we must have a pentagon at the center, and this in turn implies that we must add an additional “house” to vertex $v_{1}$, so we arrive at Figure 28.
Figure 28. 3-neighborhood of $v_{1}$ with $r(v_{1})=(3,4,4,5).$
We remark that this is exactly part of the twisted $(3,4,5,4)$-solid. To get a contradiction, we count the number of closed walks of length 6 passing through $v_{1}$ and through $v_{2}$. The non-trivial shapes that contribute to the number of closed walks of length 6 are graphed in Figure 29.
Figure 29. Possible non-trivial shapes contributing to the number of closed walks of length 6.
One can see that $v_{1}$ and $v_{2}$ share the same count for all shapes except for the double-square, for which $v_{1}$ has 1 but $v_{2}$ has 2, a contradiction. Therefore, there is no vertex $v$ satisfying $r(v)=(3,4,4,5)$ in this sub-case, and $v_{1}$ satisfies $r(v_{1})=(3,4,5,4)$. Again by counting closed walks of length 5, one can easily show that all vertices adjacent to $v_{1}$ must be adjacent to a pentagon as well, and as a consequence, $r(v)\equiv(3,4,5,4)$. The proof is complete.

### v2

@@
 Then, we use this criterion in conjunction with a computer search to find minimal examples of various types of graphs on which cospectral but non-similar vertices exist, including minimal walk-regular non-vertex-transitive graphs, which turn out to be non-planar.
-Moreover, as our main result, we classify all finite walk-regular planar graphs, proving that such graphs must be vertex-transitive.
+Moreover, as our main result, we classify all finite 3-connected walk-regular planar graphs, proving that such graphs must be vertex-transitive.
 1. Introduction
@@
 All graphs considered in this article will be finite, simple, undirected, and connected unless otherwise specified.
-The study of cospectral vertices has gained more attention recently, (see e.g., [God12, GS17, ${}^{+}$19, EKL19, KSSW20]), especially due to their significant role in quantum information theory.
+The study of cospectral vertices has gained more attention recently, (see e.g., [God12, GS17, CCT+19, EKL19, KSSW20]), especially due to their significant role in quantum information theory.
 It has been shown [God12] that for a perfect quantum state transfer to occur between two vertices in an interacting qubits network, it is necessary for these vertices to be cospectral.
@@
 Smooth analogs of walk-regular graphs were studied by Wang and the last two authors in [WWX23].
-There we show that every compact two-dimensional surface satisfying the smooth version of walk-regularity is also homogeneous—the isometry group acts transitively on the points in the surface.
+There they show that every compact two-dimensional surface satisfying the smooth version of walk-regularity is also homogeneous—the isometry group acts transitively on the points in the surface.
 The assumption that our manifold is two-dimensional is essential to the proof, and it is yet unknown if there are any higher-dimensional manifolds that are walk-regular but not homogeneous.
@@
 Theorem 1.5.
-Every walk-regular planar graph with degree $d\geq 3$ is isomorphic to the polyhedral graphs of one of the following:
+Every 3-connected walk-regular planar graph is isomorphic to the polyhedral graphs of one of the following:
   • 5 regular polyhedrons, also known as Platonic solids: tetrahedron, hexahedron, octahedron, dodecahedron, and icosahedron;
@@
 Consequently, every walk-regular planar graph is vertex-transitive.
-A direct consequence of the classical theorem of Mani [Man71] says that all planar vertex-transitive finite graphs with degree at least 3 are exactly those listed above.
-Herbert and Wilfried [FI79] employed modified Schläfli symbol to give a direct proof of this fact.
-Theorem 1.5 is a refinement of this result of Mani and Herbert–Wilfried, as walk-regularity is, in general, strictly weaker than vertex-transitivity.
+Figure 1. All the 13 semi-regular polyhedrons. The graphics are due to Wikipedia user Cyp [Cyp12], used with permission.
+A direct consequence of the classical theorem of Mani [Man71] says that all 3-connected planar vertex-transitive finite graphs are exactly those listed above.
+Fleischner and Imrich [FI79] employed the modified Schläfli symbol to give a direct proof of this fact.
+Theorem 1.5 is a refinement of this result of Mani and Fleischner–Imrich, as walk-regularity is, in general, strictly weaker than vertex-transitivity.
 The modified Schläfli symbol again plays a crucial role in our proof of Theorem 1.5.
-Figure 1. All the 13 semi-regular polyhedrons. The graphics are due to Wikipedia user Cyp [Cyp12].
+To be more detailed, we first attribute the counting of short closed walks to the counting of short cycles, and by some delicate arguments, we prove that in most 3-connected walk-regular planar graphs, every short cycle without chords must be a face, further demonstrating that the modified Schläfli symbol at each vertex is identical.
+Remark 1.6.
+Although the planarity assumption might seem excessively restrictive, it is sharp in the sense that there exists a walk-regular toroidal graph that is not vertex-transitive. See Figure 4.
 Organization of the paper
@@
 In Section 3, we give a short overview of our computer search and present a few of our minimal examples.
-The code in full is available on GitHub along with a repository of minimal examples (https://github.com/ewy-man/graph-echolocation).
+The code is available in full on GitHub, along with a repository of minimal examples (https://github.com/ewy-man/graph-echolocation).
 Section 4 is dedicated to the proof of Theorem 1.5.
@@
 Wyman was supported by NSF grant DMS-2204397.
-The second listed author extends a thanks to Richard Lange for his very patient guidance through the process of setting up a GitHub repository.
+Wyman extends a thanks to Richard Lange for his very patient guidance through setting up a GitHub repository.
 2. Closed walks of bounded length determine if two vertices are cospectral
@@
 $A^{k}_{a,a}=A^{k}_{b,b}\qquad\text{ for all $k=0,1,\ldots,n-1$.}$
-We note that the number $n-1$ of counts is sharp, since the adjacency matrix $A$ of a simple path with $n$ vertices for every odd $n=2\ell+1$ (i.e., $A_{i,j}=\delta_{|i-j|=1}$) as well as the adjacency matrix $A$ of a simple path with $n$ vertices and a self-loop attached to the last vertex for every even $n=2\ell$ (i.e., $A_{i,j}=\delta_{|i-j|=1}+\delta_{i=j=n}$) satisfies
+We note that the number $n-1$ of counts is sharp since the adjacency matrix $A$ of a simple path with $n$ vertices for every odd $n=2\ell+1$ (i.e., $A_{i,j}=\delta_{|i-j|=1}$) as well as the adjacency matrix $A$ of a simple path with $n$ vertices and a self-loop attached to the last vertex for every even $n=2\ell$ (i.e., $A_{i,j}=\delta_{|i-j|=1}+\delta_{i=j=n}$) satisfies
 $A^{k}_{\ell,\ell}=A^{k}_{\ell+1,\ell+1}\quad\text{for all}~{}k=0,1,\ldots,n-2\quad\text{but}\quad A^{n-1}_{\ell,\ell}<A^{n-1}_{\ell+1,\ell+1}.$
@@
 An exhaustive list of minimal examples of graphs with cospectral, non-similar vertices under regularity conditions of varying strength can be found on GitHub (https://github.com/ewy-man/graph-echolocation) along with the code used to generate them.
-The code is written in the Julia language (see [BEKS17, ${}^{+}$21]) and contains a number of tools:
+The code is written in the Julia language (see [BEKS17, FBS+21]) and contains a number of tools:
   1. (1)
@@
   4. (4)
-    A number of other convenience functions, such as those that determine if a graph is walk-regular, or if there exist non-similar cospectral pairs.
+    A number of other convenience functions, such as those that determine if a graph is walk-regular or if there exist non-similar cospectral pairs.
 Once loaded into the Julia REPL, these functions can be called to generate the minimal examples below and to check that they are indeed minimal.
@@
 3.3. Minimal walk-regular, non-vertex-transitive graphs
-There are no walk-regular, non-vertex-transitive graphs which have less than 12 vertices.
+There are no walk-regular, non-vertex-transitive graphs, which have less than 12 vertices.
 There are exactly four minimal examples with twelve vertices, one each of degree 4, 5, 6, and 7.
@@
 The degree $4$ and $5$ examples are depicted in Figure 3.
+Furthermore, even though the degree 4 example is not planar, it can be embedded into a torus, as shown in Figure 4.
 Refer to caption      Refer to caption
 Figure 3. These graphs and their edge-complements are the only walk-regular, non-vertex-transitive graphs on $12$ vertices. Red vertices belong to one orbit and blue vertices belong to the other.
+Figure 4. A walk-regular non-vertex-transitive toroidal graph.
 4. Walk-regular planar graphs
-Here we will set up the tools we will need in order to classify the planar walk-regular graphs and prove Theorem 1.5.
+4.1. Auxiliary lemmas
+Here we set up the tools we will need in order to classify 3-connected walk-regular planar graphs and prove Theorem 1.5.
+The first tool shall be used to obtain the number of short closed walks through a vertex from the number of cycles through it.
+Let $G$ be a $d$-regular graph with $d\geq 3$.
+A closed walk is said to be contractible if it traces no cycles.
+For a vertex $v$ and $m\geq 3$, we let
+$\mathcal{C}_{m}(v)=\#\{\text{$m$-cycles through $v$}\}.$
+As every closed walk of length 3 traces a triangle in one of two directions,
+$(A^{3})_{v,v}=2\mathcal{C}_{3}(v).$
+Starting at $v$ there are exactly $d(d-1)+d^{2}$ contractible closed walks with length 4. Adding in the ones tracing 4-cycles, we have
+$(A^{4})_{v,v}=2d^{2}-d+2\mathcal{C}_{4}(v).$
+Now we assume that $\mathcal{C}_{3}(v_{i})\equiv\mathcal{C}_{3}$ for all vertices $v_{i}$.
+Note that every closed walk $\overline{va_{1}a_{2}a_{3}a_{4}v}$ traces either a 5-cycle or an edge attached to a triangle; among the latter ones, there are exactly $2d\mathcal{C}_{3}$ with $v=a_{2}$, $2d\mathcal{C}_{3}$ with $v=a_{3}$, $2(d-2)\mathcal{C}_{3}$ with $a_{1}=a_{4}$ and $v\notin\{a_{2},a_{3}\}$, $2(d-2)\mathcal{C}_{3}$ with $a_{1}=a_{3}$ and $a_{2}\notin\{v,a_{3}\}$, $2(d-2)\mathcal{C}_{3}$ with $a_{2}=a_{4}$ and $a_{3}\notin\{v,a_{1}\}$, and $2\mathcal{C}_{3}$ with $a_{1}=a_{3}$ and $a_{2}=a_{4}$.
+In total we have
+$(A^{5})_{v,v}=10(d-1)\mathcal{C}_{3}+2\mathcal{C}_{5}(v).$
+Furthermore, assume that $\mathcal{C}_{4}(v_{i})\equiv\mathcal{C}_{4}$ for all vertices $v_{i}$.
+If $G$ is triangle-free (i.e., $\mathcal{C}_{3}=0$), similarly, we count the closed walks starting at $v$ with length 6: $d^{3}+2d^{2}(d-1)+2d(d-1)^{2}$ of them are contractible, $2\cdot 2d\mathcal{C}_{4}+4\cdot 2(d-2)\mathcal{C}_{4}+2\cdot 2\mathcal{C}_{4}$ of them trace some 4-cycles with attached edges, and the remaining $2\mathcal{C}_{6}(v)$ trace some 6-cycles.
+Therefore,
+$(A^{6})_{v,v}=5d^{3}-6d^{2}+2d+12(d-1)\mathcal{C}_{4}+2\mathcal{C}_{6}(v).$
+If $\mathcal{C}_{3}=1$, compared to the triangle-free case, there are two fewer contractible closed walks of length 6 but four more such walks that trace the triangle through $v$ twice.
+In this case, it follows that
+$(A^{6})_{v,v}=5d^{3}-6d^{2}+2d+12(d-1)\mathcal{C}_{4}+2\mathcal{C}_{6}(v)+2.$
+A direct consequence of the above formulas is that
+Lemma 4.1.
+Let $G$ be a walk-regular graph. Then
+  1. (1)
+    $\mathcal{C}_{3}(v)\equiv\mathcal{C}_{3},\,\mathcal{C}_{4}(v)\equiv\mathcal{C}_{4}$, and $\mathcal{C}_{5}(v)\equiv\mathcal{C}_{5}$ for all vertices $v_{i}$;
+  2. (2)
+    if $\mathcal{C}_{3}\leq 1$, $\mathcal{C}_{6}(v)$ is also constant over all vertices $v_{i}$.
 […]
@@
 $f=\sum_{v}p(v),\qquad\text{where}\ \ p(v):=\sum_{i=1}^{d}\frac{1}{r_{i}(v)}.$
-Using Euler’s formula $n+f-e=2$ and $e=nd/2$, we get
-(4.1) $\frac{1}{n}\sum_{v}p(v)=\frac{d-2}{2}+\frac{2}{n}$
-We let $k(v)=\#\{1\leq i\leq d:r_{i}(v)=r_{1}(v)\}$. An easy observation is
+Using Euler’s identity $n+f-e=2$ and $e=nd/2$, we get
+(4.1) $\frac{1}{n}\sum_{v}p(v)=\frac{d-2}{2}+\frac{2}{n}.$
+In order to prove Theorem 1.5, we shall prove the following key lemma.
 Lemma 4.2.
-In a walk-regular planar graph with $d\geq 3$,
-[…]
-Proof.
-First we let $r_{1}=\min_{v}r_{1}(v)$. Then it follows from (4.1) that
-$\frac{d}{r_{1}}\geq\frac{d-2}{2}+\frac{2}{n}.$
-Thus $r_{1}\in\{3,4,5\}$ for $d=3$, and $r_{1}=3$ for $d=4$ or $5$.
-Let $v_{1}$ be a vertex such that $r_{1}(v_{1})=r_{1}$.
-Suppose that there is a vertex $v_{2}$ that satisfies $r_{1}(v_{2})>r_{1}$.
-If $r_{1}\in\{3,5\}$, we have
-$A^{r_{1}}_{v_{2},v_{2}}=0<2k(v_{1})=A^{r_{1}}_{v_{1},v_{1}}.$
-If $r_{1}=4$, then $d=3$ and
-$A^{4}_{v_{2},v_{2}}=15<15+2k(v_{1})=A^{4}_{v_{1},v_{1}}.$
-In any case, $v_{1}$ and $v_{2}$ are not cospectral. Hence $r_{1}(v)\equiv r_{1}$ provided that the graph is walk-regular. By a similar closed walk counting argument, we conclude that
-$A_{v,v}^{r_{1}}-A_{v^{\prime},v^{\prime}}^{r_{1}}=2k(v)-2k(v^{\prime})$
-for any pair of vertices $v,v^{\prime}$, and thus $k(v)\equiv k$ by walk-regularity. ∎
-In order to prove Theorem 1.5, we shall prove the following lemma.
-Lemma 4.2.
-For any walk-regular planar graph with degree $d\geq 3$, the modified Schläfli symbol is constant over all vertices, that is
+For any 3-connected walk-regular planar graph with degree $d\geq 3$, the modified Schläfli symbol is constant over all vertices, that is
 (4.2) $r(v)\equiv(r_{1},r_{2},\cdots,r_{d})\qquad\textit{for all vertices}~{}v.$
-However, the proof of this lemma appears to be quite delicate.
-The 3-regular planar graph, as illustrated in Figure 5, features cospectral vertices $v_{1}$ and $v_{2}$ with $r(v_{1})=(3,5,5)$ but $r(v_{2})=(3,5,6)$.
-Despite vertex $v_{1}$ having two neighboring pentagons and $v_{2}$ only one, the adjacent triangle $\{v_{5},v_{7},v_{9}\}$ contributes two additional closed walks of length 5 to $v_{2}$.
-Moreover, this graph is ‘almost’ walk-regular, with all vertices except $v_{4}$ being cospectral.
-Therefore, establishing (4.2) succinctly appears challenging.
-Instead, we justify (4.2) in each individual case.
+However, establishing (4.2) succinctly appears challenging. Instead, we will justify (4.2) in each individual case.
 We remark that, once we prove Lemma 4.2, it is routine to check that Theorem 1.5 holds in each case.
 Indeed, the numbers of vertices, edges, and faces are determined by Lemma 4.2 and Euler’s identity.
-The graph can then be reconstructed in a deterministic fashion, starting with an arbitrary polygon as the outer polygon.
+The graph can then be reconstructed in a deterministic fashion, starting with an arbitrary polygon as the outer face.
 We illustrate this fact for the most complicated case when $r(v)\equiv(3,4,5,4)$ in Figure 5.
@@
 Figure 5. Step by step reconstruction of the $(3,4,5,4)$-solid.
-Figure 5. A 3-regular planar graph on 10 vertices.
-Let $v_{1}$ be a vertex such that $p(v_{1})=\max_{v}p(v)$.
-In view of Lemma 4.1, to prove Lemma 4.2, we only need to discuss the sixteen cases enumerated below.
+We fix a vertex $v_{1}$ such that $p(v_{1})=\max_{v}p(v)$. Let
+$r_{1}=r_{1}(v_{1})\quad\text{and}\quad k=\#\{1\leq i\leq d:r_{i}(v_{1})=r_{1}(v_{1})\}.$
+In view of (4.1), to prove Lemma 4.2, we only need to discuss the fifteen cases enumerated below.
 For the convenience of the reader, we have also included an exhaustive list of the possible walk-regular planar graphs belonging to each case.
@@
 (15) 5, 3, 4 $(3,3,3,3,4)$- and $(3,3,3,3,5)$-solids.
-(16) 5, 3, $\leq$ 3
-Of these cases, (2) and (8) are simply impossible to realize. It is also impossible to realize case (16), since
-$p(v)\leq\frac{1}{3}+\frac{1}{3}+\frac{1}{3}+\frac{1}{4}+\frac{1}{4}=\frac{3}{2}$
-violates (4.1). One quickly sees the symbols $r(v)$ of the graphs realizing cases (1), (4), (7), (10), and (14) must be constant.
-Next, we focus on cases (5), (11), (9), and (15), which are easier to resolve. The remaining cases (3), (6), (12), and (13) are more involved. We will treat these cases after.
+Table 1. 15 possible cases.
+To effectively use Lemma 4.1 in discussions, we need another tool to convert the number of short cycles into the one of faces.
+We label an $m$-cycle $C=\overline{a_{1}a_{2}\cdots a_{m}a_{1}}$ by its inner degree sequence
+$d_{*}(C)=(d_{*}(a_{1}),d_{*}(a_{2}),\cdots,d_{*}(a_{m})),$
+where $d_{*}(a_{i})$ is the number of incident edges of $a_{i}$ inside $C$; here an object (vertex, edge, cycle, face, etc) is said to be “inside $C$” if it lies in the region bounded by $C$ in the planar embedding.
+Again, $d_{*}(C)$ is uniquely defined, once we choose $a_{1}$ and the direction to minimize the sequence lexicographically.
+Clearly, $C$ is a face if and only if $d_{*}(C)=(0,0,\cdots,0)$ or $(d-2,d-2,\cdots,d-2)$ (the latter means that $C$ is the unique outer face of the embedding).
+For $m=3$, we have
+Lemma 4.3.
+In a 3-connected walk-regular planar graph, every triangle is a face.
+Note that a cycle $C$ in $G$ with length $\geq 4$ is not necessarily a face.
+The typical situation is that $C$ has a chord, an edge not in $C$ whose endpoints lie in $C$.
+For instance, the (3,3,3,3,4)-solid contains a 4-cycle with one chord, a 5-cycle with one chord and a 5-cycle with two chords; none of them are faces.
+We also observe that the 4-cycle encircling four triangles in the octahedron, the 5-cycle encircling five triangles in the icosahedron, and the 5-cycle encircling three triangles and a quadrangular face in the 4-gonal antiprism are not faces, even though all of them are chordless.
+Indeed, by excluding these cases, we have the next two lemmas for 4- and 5-cycles.
+Lemma 4.4.
+In a 3-connected walk-regular planar graph which is not isomorphic to the octahedron, every chordless 4-cycle is a face.
+Lemma 4.5.
+In a 3-connected walk-regular planar graph which is not isomorphic to the 4-gonal antiprism or the icosahedron, every chordless 5-cycle is a face.
+Next, we will provide the proofs of Lemmas 4.34.5 separately in the cases $d=3,4,5$, and conclude the proof of Lemma 4.2, and thus Theorem 1.5 for each case.
+4.2. The cubic cases
+Assume that $G$ is cubic (i.e., $3$-regular).
+When a chordless $m$-cycle $C$ in $G$ is not a face, the $3$-connectedness yields that $\geq 3$ vertices on $C$ have incident edges inside $C$, and $\geq 3$ of them have incident edges outside $C$, hence $m\geq 6$.
+This justifies Lemmas 4.34.5 for all cubic cases.
+Consequently, for any vertex $v$, we have
+$r_{1}(v)\equiv r_{1}\quad\text{and}\quad k(v):=\#\{1\leq i\leq 3:r_{i}(v)=r_{1}\}\equiv k.$
+Among the cases (1)–(9) in Table 1, (2) and (8) are simply impossible to realize.
+Also, one quickly sees the symbols $r(v)$ of the graphs realizing cases (1), (4), and (7) must be constant, where the corresponding determined solids are the tetrahedron, the hexahedron, and the dodecahedron.
+In case (5) ($d=3,r_{1}=4,k=2$), the face $f_{3}(v_{1})$ surrounded by $m=r_{3}(v_{1})\geq 5$ quadrangular faces determines the unique pattern up to isomorphism.
+Therefore, $r(v)\equiv(4,4,m)$ and the graph is isomorphic to the $m$-gonal prism.
+For the other three cases (with $d=3,k=1$), we need to consider the second shortest cycles in $G$, that is, the cycles with length
+$\ell=\min\{m>r_{1}:\mathcal{C}_{m}(v)>0\text{ for some vertex }v\}.$
+It is clear that $\ell\leq r_{2}(v_{1})$. We will claim in each individual case that
+Lemma 4.6.
+In a 3-connected walk-regular planar graph with $d=3$ and $k=1$, every $\ell$-cycle is a face.
 Case (3)
-($d=3,r_{1}=4,k=2$) For each $m=r_{3}(v_{1})\geq 5$, the face $f_{3}(v_{1})$ surrounded by $m$ quadrangular faces determines the unique pattern up to isomorphism.
-In this case, $r(v)\equiv(4,4,m)$ and the graph is isomorphic to the $m$-gonal prism.
+($d=3,r_{1}=3,k=1$) Note that each $\ell$-cycle is chordless, and it should be adjacent to exactly $\ell/2$ triangles, and thus $\ell$ must be even.
+Suppose that an $\ell$-cycle $C$ is not a face.
+Then by the 3-connectedness, $C$ has $\geq 3$ neighbors inside $C$ and $\geq 3$ neighbors outside $C$.
+Since $C$ is chordless and each vertex must be adjacent to a triangle, we must have $\ell\geq 12$.
+However, in this situation one has $r_{2}(v_{1})\geq 12$ and
+$p(v_{1})\leq\frac{1}{3}+\frac{1}{12}+\frac{1}{12}=\frac{1}{2},$
+which violates (4.1). This proves Lemma 4.6 for $r_{1}=3$, and the possible $\ell=4,6,8$ or $10$.
+Assume $\ell=4$ or $6$.
+It follows from Lemma 4.1 that $\mathcal{C}_{4}(v)$ or $\mathcal{C}_{6}(v)$ is constant for all vertices $v$.
+For $\ell=8$, we can count the number of $8$-cycles though a vertex in its 4-neighborhood to obtain
+$(A^{8})_{v,v}=591+2\mathcal{C}_{8}(v).$
+For $\ell=10$, a similar counting in the 5-neighborhood of the vertex $v$ provides
+$(A^{10})_{v,v}=4223+2\mathcal{C}_{10}(v).$
+Hence whatever $\ell$ is, we have $\mathcal{C}_{\ell}(v)\equiv\mathcal{C}_{\ell}$.
+Since 3 is odd, drawing a triangular face on which each vertex is adjacent to exactly one face with length $\ell$ is impossible.
+The only other possibility is $\mathcal{C}_{\ell}=2$, and it follows that $r(v)\equiv(3,\ell,\ell)$.
+The corresponding solids are the triangular prism, the $(3,6,6)$-solid, the $(3,8,8)$-solid and the $(3,10,10)$-solid respectively.
 Case (6)
-($d=4,r_{1}=3,k=3$) For each $m=r_{4}(v_{1})\geq 4$, the face $f_{4}(v_{1})$ surrounded by $m$ triangular faces determines the unique pattern up to isomorphism.
-In this case, $r(v)\equiv(3,3,3,m)$ and the graph is isomorphic to the $m$-gonal antiprism.
-Starting from the next case, we often need to count the closed walks of length $m$ for some $m\geq 4$ and two adjacent vertices $v,v^{\prime}$ with $r(v)=r(v_{1})\neq r(v^{\prime})$.
-Our strategy of counting is to consider only the $\lfloor\frac{m}{2}\rfloor$-neighborhood of $\{v,v^{\prime}\}$, which turns out to be determined by the values of $d,r_{1},k$ and $r(v_{1})$, with at most two possibilities.
+($d=3,r_{1}=4,k=1$) We first claim that every cycle has even length.
+If this was false, we let $\ell^{\prime}=\min\{\text{odd }m:\mathcal{C}_{m}(v)>0\text{ for some vertex }v\}$, and let $C$ be a minimal $\ell^{\prime}$-cycle, in the sense that no $\ell^{\prime}$-cycles are properly contained in the closed region bounded by $C$.
+Suppose that a vertex $b_{1}$ is inside $C$.
+The walk-regularity implies that, starting at $b_{1}$ there is at least one closed walk with length $\ell^{\prime}$, denoted by $C^{\prime}=\overline{b_{1}b_{2}\cdots b_{\ell^{\prime}}b_{1}}$.
+By the minimality of $\ell^{\prime}$ and $C$, $C^{\prime}$ should be a cycle, and some of vertices on $C^{\prime}$ lie outside $C$.
+Let $b_{i},b_{j}$ be the first and the last intersection vertices of $C$ and $C^{\prime}$ respectively, that is, $i=\min\{m:b_{m}~{}\text{lies on}~{}C\}$ and $j=\max\{m:b_{m}~{}\text{lies on}~{}C\}$.
+Such $b_{i}$ and $b_{j}$ divide $C$ into two paths $\mathcal{P}_{1},\mathcal{P}_{2}$ with length $\ell_{1},\ell_{2}$, and since $\ell_{1}+\ell_{2}=\ell^{\prime}$ is odd, one of $\ell_{1},\ell_{2}$ (say $\ell_{1}$) has the same parity as $\ell^{\prime}+i-j$.
+As the union of two paths $\overline{b_{i}b_{i-1}\cdots b_{1}b_{\ell^{\prime}}b_{\ell^{\prime}-1}\cdots b_{j}}$ and $\mathcal{P}_{2}$ form a cycle with odd length $\ell^{\prime}+i-j+\ell_{2}$ inside the closed region bounded by $C$, it follows again from the minimality of $\ell^{\prime}$ and $C$ that $\ell^{\prime}+i-j+\ell_{2}>\ell^{\prime}$.
+But now, one has the union of two paths $\overline{b_{i}b_{i+1}\cdots b_{j}}$ and $\mathcal{P}_{1}$ as a closed walk with odd length $j-i+\ell_{1}<\ell^{\prime}$, which still violates the minimality of $\ell^{\prime}$.
+Therefore, $C$ has no inner neighbors, and it should be an inner face in the planar embedding; however, this is also impossible, since at least one adjacent 4-cycle of $C$ shares three vertices with $C$, of which the middle one must have an inner incident edge contradicting Lemma 4.4.
+The claim is justified.
+Using (4.1), we obtain $\ell\leq r_{2}(v_{1})<8$.
+By the claim, the only possibility is $\ell=r_{2}(v_{1})=6$.
+For a 6-cycle $C$, it has either two or three adjacent $4$-cycles.
+In the former situation, $C$ has only two inner neighbors and two outer neighbors, which violates 3-connectedness.
+For the latter one, to be 3-connected all of the three adjacent $4$-cycles must lie outside $C$ whenever $C$ is not the outer face.
+This proves that every $6$-cycle is a face, and justifies Lemma 4.6 for $r_{1}=4$.
+In view of Lemma 4.1, $\mathcal{C}_{6}(v)\equiv\mathcal{C}_{6}=1$ or $2$. For $\mathcal{C}_{6}=2$, we have $r(v)\equiv(4,6,6)$, and the graph is isomorphic to the $(4,6,6)$-solid.
+Now we consider the sub-case $\mathcal{C}_{6}=1$.
+It follows from (4.1) and the claim that $r_{3}(v_{1})=8$ or $10$.
+For a vertex $v$ with $r(v)=(4,6,m)$, by counting in the 4- and 5-neighborhood of $v$ we obtain
+$(A^{8})_{v,v}=809+\begin{cases}\,2\quad\text{if }m=r_{3}(v_{1})=8,\\ \,0\quad\text{if }m>r_{3}(v_{1})=8,\end{cases}$
+and
+$(A^{10})_{v,v}=6063+\begin{cases}\,2\quad\text{if }m=r_{3}(v_{1})=10,\\ \,0\quad\text{if }m>r_{3}(v_{1})=10.\end{cases}$
+Hence, $r(v)\equiv(4,6,8)$ or $\equiv(4,6,10)$, which corresponds to the $(4,6,8)$-solid or the $(4,6,10)$-solid respectively.
 Case (9)
-($d=3,r_{1}=5,k=1$) Each face with $m\geq 6$ edges must be surrounded by exact $m/2$ pentagonal faces and thus $m$ is even.
-As $p(v_{1})>\frac{1}{2}$ by (4.1), the only possibility is $r(v_{1})=(5,6,6)$.
-Suppose $r(v)\not\equiv(5,6,6)$.
-Then there exists a pentagonal face with two adjacent vertices $v,v^{\prime}$ such that $r(v)=(5,6,6)$ and $r(v^{\prime})=(5,6,m)$ with $m>6$.
-The $3$-neighborhood of $\{v,v^{\prime}\}$ is unique, with 23 vertices (Figure 6).
-This gives
-$A_{v,v}^{6}=91>89=A^{6}_{v^{\prime},v^{\prime}}$
-Hence $r(v)\equiv(5,6,6)$, the graph is isomorphic to the $(5,6,6)$-solid.
-Figure 6. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(5,6,6)$ and $r(v^{\prime})=(5,6,m)$ with $m>6$.
+($d=3,r_{1}=5,k=1$) Using (4.1) we immediately get $\ell=r_{2}(v_{1})=6$. We will prove that every $6$-cycle is a face in this case, and conclude the proof of Lemma 4.6.
+Proof of Lemma 4.6 for $r_{1}=5$.
+Suppose that a 6-cycle $C=\overline{a_{1}a_{2}\cdots a_{6}a_{1}}$ is not a face.
+Without loss of generality, we further assume that such $C$ is minimal among all the 6-cycles that are not faces.
+More precisely, this minimality of $C$ implies that any 6-cycle properly contained in the closed region bounded by $C$ should be a face.
+By the 3-connectedness of the graph, $C$ has exactly three inner incident edges and three outer incident edges.
+There is a unique (up to isomorphism) possible neighborhood of $C$ such that $\mathcal{C}_{5}(a_{i})=1$ for all vertices $a_{i}$ on $C$, as depicted in Figure 6, where $C$ is adjacent to a face $\overline{b_{2}a_{4}a_{5}a_{6}b_{3}b_{2}}$ with length 5 inside $C$ and a pentagon face outside $C$ bounded by the path $\overline{a_{1}a_{2}a_{3}}$.
+Note that the inner neighbor $b_{1}$ of $a_{2}$ is not adjacent to $b_{2}$ or $b_{3}$, and hence, both the face bounded by the path $\overline{b_{1}a_{2}a_{1}a_{6}b_{3}}$ and the one bounded by $\overline{b_{1}a_{2}a_{3}a_{4}b_{2}}$ have length $>5$.
+Besides these two faces, the remaining adjacent face of $b_{1}$ should be of length $5$ so that $\mathcal{C}_{5}(b_{1})=1$.
+However, no matter which adjacent face of $b_{1}$ is of length $6$, the minimality of $C$ implies that $\mathcal{C}_{6}(b_{1})<\mathcal{C}_{6}(a_{2})$, which contradicts Lemma 4.1.
+Figure 6. An impossible 6-cycle with $r_{1}=5$.
+Now using Lemmas 4.1 and 4.6 we have $\mathcal{C}_{6}(v)\equiv\mathcal{C}_{6}=1$ or $2$.
+A pentagon face on which each vertex is adjacent to exactly one face with length 6 is simply impossible, and hence $\mathcal{C}_{6}=2$.
+This yields that $r(v)\equiv(5,6,6)$, and the graph is isomorphic to the $(5,6,6)$-solid.
+4.3. The 4-regular cases
+Now we assume that $G$ is 4-regular. We first justify that every triangle in $G$ is a face:
+Proof of Lemma 4.3 for $d=4$.
+Suppose that a triangle $C$ is not a face.
+By the 3-connectedness, each vertex $a_{i}$ has at least one incident edge inside $C$ and at least one incident edge outside $C$.
+Therefore, it must have $d_{*}(C)=(1,1,1)$.
+However, this is simply impossible as the total degree of the induced subgraph of the vertices inside $C$ is odd.
+∎
+Consequently, the octahedron is determined as $r(v)\equiv(3,3,3,3)$ in case (10).
+For case (11) ($d=4,r_{1}=3,k=3$), the face $f_{4}(v_{1})$ surrounded by $m=r_{4}(v_{1})\geq 4$ triangular faces determines the unique pattern up to isomorphism, hence $r(v)\equiv(3,3,3,m)$ and the graph is isomorphic to the $m$-gonal antiprism.
+The other two cases (12) and (13) are more involved. We need to prove that every 4-cycle is either a face or the boundary of two adjacent triangular faces there.
+Proof of Lemma 4.4 for $d=4$.
+Suppose that a chordless 4-cycle $C=\overline{a_{1}a_{2}a_{3}a_{4}a_{1}}$ is not a face.
+By the 3-connectedness, at least three of $\{a_{1},a_{2},a_{3},a_{4}\}$ have incident edges inside $C$, and at least three of them have incident edges outside $C$.
+Without loss of generality, we assume that such $C$ is minimal, that is, any chordless 4-cycle properly contained in the closed region bounded by $C$ should be a face.
+Since the total degree of the subgraph inside $C$ is even, the number of incident edges of $\{a_{1},a_{2},a_{3},a_{4}\}$ inside $C$ equals $4$.
+The possible inner degree sequence of $C$ is $(0,1,1,2),(0,1,2,1)$ or $(1,1,1,1)$.
+The case $d_{*}(C)=(0,1,2,1)$ is simply ruled out, as $\{a_{2},a_{4}\}$ should not be a cut-set of the graph.
+Case: $d_{*}(C)=(0,1,1,2)$.
+Note that $\mathcal{C}_{3}=\mathcal{C}_{3}(v_{1})\geq 1$.
+In this case, $\mathcal{C}_{3}\geq 2$ is impossible since $\{a_{1},a_{2},a_{3},a_{4}\}$ has $\geq 3$ neighbors inside $C$ and $\geq 3$ neighbors outside $C$ (by the 3-connectedness).
+It follows that $\mathcal{C}_{3}=1$, and every 4-cycle inside $C$ is chordless so is a face.
+Using (4.1), we have $\mathcal{C}_{4}=\mathcal{C}_{4}(v_{1})\geq 2$.
+Lemma 4.3 implies that every vertex inside $C$ is surrounded by at most one face with length $>4$.
+Let $b_{1},b_{2}$ be the two neighbors of $a_{4}$ inside $C$.
+When $\overline{b_{1}b_{2}}$ is an edge, to have $\mathcal{C}_{3}(a_{3})=\mathcal{C}_{3}(a_{4})=1$, $a_{2}$ and $a_{3}$ have exactly one common neighbor $u$, and this gives a cut-set $\{a_{4},u\}$ if $u$ lies inside $C$ or a cut-set $\{a_{1},u\}$ if $u$ lies outside $C$, which violates the 3-connectedness.
+Now the only possibility left is that one of $b_{1},b_{2}$ (say $b_{1}$) is adjacent to $a_{3}$, and one of the neighbors of $a_{1}$ outside $C$ is adjacent to $a_{2}$.
+Let $b_{3}$ be the neighbor of $a_{2}$ inside $C$.
+Since the face bounded by $\overline{b_{2}a_{4}a_{1}a_{2}b_{3}}$ has length $>4$, we have $\mathcal{C}_{4}=2$, and both the face bounded by $\overline{b_{2}a_{4}b_{1}}$ and the one bounded by $\overline{b_{3}a_{2}a_{3}b_{1}}$ have length $4$, as drawn in Figure 7.
+Then, to have $\mathcal{C}_{3}(b_{1})+\mathcal{C}_{4}(b_{1})=3$, the undetermined face with $\overline{b_{1}b_{3}}$ should have length $>4$, contradicting the fact that $\mathcal{C}_{3}(b_{3})+\mathcal{C}_{4}(b_{3})=3$.
+Figure 7. An impossible 4-cycle with $d_{*}(C)=(0,1,1,2)$.
+Case: $d_{*}(C)=(1,1,1,1)$.
+One special situation is that $\{a_{1},a_{2},a_{3},a_{4}\}$ has a common neighbor inside $C$, in which $\mathcal{C}_{3}=4$ and the graph is isomorphic to the octahedron.
+If this is not the case, again by the 3-connectedness, there are at least three distinct neighbors of $\{a_{1},a_{2},a_{3},a_{4}\}$ inside $C$ as well as at least three distinct ones outside $C$.
+From this we know that $\mathcal{C}_{3}=1$, one edge (say $\overline{a_{1}a_{4}}$) of $C$ lies on an incident triangle inside $C$, and the opposite edge ($\overline{a_{2}a_{3}}$) of $C$ lies on an incident triangle outside $C$.
+Denote by $b_{1},b_{2},b_{3}$ the neighbors of $a_{1},a_{2},a_{3}$ inside $C$ respectively.
+Note that there are no chorded $4$-cycles inside $C$ or else $\mathcal{C}_{3}\geq 2$, and hence every $4$-cycle inside $C$ is a face by the minimality of $C$, thus $\mathcal{C}_{4}\leq 3$.
+If $\mathcal{C}_{4}=3$, then none of the faces inside $C$ has length $>4$, and it forces $\overline{b_{1}b_{2}b_{3}b_{1}}$ to be a triangle but not a face, which violates Lemma 4.3.
+Now we suppose $\mathcal{C}_{4}=2$.
+Once $b_{2},b_{3}$ are adjacent, to have $\mathcal{C}_{4}(a_{2})=\mathcal{C}_{4}(a_{3})=2$, both the face bounded by $\overline{b_{2}a_{2}a_{1}b_{1}}$ and the one bounded by $\overline{b_{3}a_{3}a_{4}b_{1}}$ have length $>4$, which yields that $\mathcal{C}_{4}(b_{1})\leq 1$.
+Hence any graph with $\overline{b_{2}b_{3}}$ is ruled out.
+Without $\overline{b_{2}b_{3}}$, the face bounded by $\overline{b_{2}a_{2}a_{3}b_{3}}$ has length $>4$.
+To have $\mathcal{C}_{4}(b_{2})=\mathcal{C}_{4}(b_{3})=2$, both $b_{2},b_{3}$ must be adjacent to $b_{1}$, which gives a cut-set $\{b_{2},b_{3}\}$ of the graph (Figure 8), a contradiction.
+Figure 8. An impossible 4-cycle with $d_{*}(C)=(1,1,1,1)$.
 Case (12)
-($d=5,r_{1}=3,k=4$) As $p(v_{1})>\frac{3}{2}$ by (4.1), $r_{5}(v_{1})=4$ or $5$.
-When $r_{5}(v_{1})=4$, we suppose $r_{5}(v)\not\equiv 4$.
-Then there exist two adjacent vertices $v,v^{\prime}$ such that $r_{5}(v)=4<r_{5}(v^{\prime})$.
-Counting closed walks of length $4$ in the $2$-neighborhood of $\{v,v^{\prime}\}$ (Figure 7), we have
-$A_{v,v}^{4}=59>57=A_{v^{\prime},v^{\prime}}^{4}.$
-Hence $r(v)\equiv(3,3,3,3,4)$, the graph is isomorphic to the $(3,3,3,3,4)$-solid.
-Figure 7. Two possible patterns of the $2$-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,3,3,3,4)$ and $r(v^{\prime})=(3,3,3,3,m)$ with $m>4$.
-In the case that $r_{5}(v_{1})=5$, $r_{5}(v)\not\equiv 5$, there must exist two adjacent vertices $v,v^{\prime}$ with $r_{5}(v)=5<r_{5}(v^{\prime})$.
-Counting closed walks of length 5 in the 2-neighborhood of $\{v,v^{\prime}\}$ (Figure 8) gives
-$A_{v,v}^{5}=182>180=A_{v^{\prime},v^{\prime}}^{5}.$
-Hence $r(v)\equiv(3,3,3,3,5)$, the graph is isomorphic to the $(3,3,3,3,5)$-solid.
-Figure 8. Two possible patterns of the $2$-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,3,3,3,5)$ and $r(v^{\prime})=(3,3,3,3,m)$ with $m>5$.
-We now move on to the more involved cases.
-Case (3)
-($d=3,r_{1}=3,k=1$) Denote the vertex set of the triangular face $f_{1}(v_{1})$ by $\{v_{1},v_{2},v_{3}\}$, where $v_{2}$ does not belong to the face $f_{2}(v_{1})$.
-Note that every face with $m\geq 4$ edges should be surrounded by exact $m/2$ triangular faces, and thus $m$ must be even.
-Suppose $r_{2}(v_{1})>10$.
-Then
-$p(v_{1})\leq\frac{1}{3}+\frac{1}{12}+\frac{1}{12}=\frac{1}{2},$
-which violates (4.1). Hence $r_{2}(v_{1})=4,6,8$ or $10$.
-Sub-case: $r_{2}(v_{1})=4$
-If $r_{3}(v_{1})>4$, it follows that $r_{2}(v_{2})>4$ and
-$A_{v_{1},v_{1}}^{4}=17>15=A_{v_{2},v_{2}}^{4}.$
-Hence the walk-regularity implies $r_{3}(v_{1})=4$. In this case, the only possibility is $r(v)\equiv(3,4,4)$, and the graph is isomorphic to the triangular prism.
-Sub-case: $r_{2}(v_{1})=6$
-Suppose $r_{3}(v_{1})>6$.
-Then $r_{2}(v_{2})>6$, otherwise we have $r(v_{3})=(3,6,6)$ that violates the minimality of $p(v_{1})$.
-Counting the closed walks with length $6$ in the $3$-neighborhood of $\{v_{1},v_{2}\}$ (Figure 9)
-$A_{v_{1},v_{1}}^{6}=91>89=A_{v_{2},v_{2}}^{6}.$
-Thus $r(v_{1})=(3,6,6)$ by the walk-regularity.
-Figure 9. 3-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,6,m)$ with $m>6$.
-Now we claim that $r(v)\equiv(3,6,6)$.
-Indeed, whenever this is false, there exists a triangular face with vertex set $\{v,v^{\prime},v^{\prime\prime}\}$ such that $r(v)=(3,6,6)$ and $r(v^{\prime})=(3,6,m)$ with $m>6$ (Figure 10), which gives
-$A_{v,v}^{6}=93>91=A_{v^{\prime},v^{\prime}}^{6}.$
-Hence $r(v)\equiv(3,6,6)$, and the graph is isomorphic to the $(3,6,6)$-solid.
-Figure 10. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,6,6)$ and $r(v^{\prime})=(3,6,m)$ with $m>6$.
-Sub-case: $r_{2}(v_{1})=8$
-As in the previous sub-case, $r_{3}(v_{1})>8$ implies $r_{2}(v_{2})>8$, and a counting in the $4$-neighborhood of the pair $\{v_{1},v_{2}\}$ (Figure 11) provides
-$A_{v_{1},v_{1}}^{8}=593>591=A_{v_{2},v_{2}}^{8},$
-thus $r(v_{1})=(3,8,8)$.
-Figure 11. 4-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,8,m)$ with $m>8$.
-If $r(v)\equiv(3,8,8)$ is violated, there is a triangular face with vertices $v,v^{\prime}$ such that $r(v)=(3,8,8)$ and $r(v^{\prime})=(3,8,m)$ with $m>8$ (Figure 12), and
-$A_{v,v}^{8}=595>593=A_{v^{\prime},v^{\prime}}^{8}.$
-This shows that $r(v)\equiv(3,8,8)$, and the graph is isomorphic to the $(3,8,8)$-solid.
-Figure 12. 4-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,8,8)$ and $r(v^{\prime})=(3,8,m)$ with $m>8$.
-Sub-case: $r_{2}(v_{1})=10$
-Again, $r_{3}(v_{1})>10$ implies $r_{2}(v_{2})>10$, and a counting in the 5-neighborhood of $\{v_{1},v_{2}\}$ (Figure 13) gives
-$A_{v_{1},v_{1}}^{10}=4225>4223=A_{v_{2},v_{2}}^{10},$
-thus $r(v_{1})=(3,10,10)$.
-Figure 13. 5-neighborhood of $\{v_{1},v_{2}\}$, where $r(v_{1})=(3,10,m)$ with $m>10$.
-If $r(v)\equiv(3,10,10)$ is violated, there is a triangular face with vertices $v,v^{\prime}$ such that $r(v)=(3,10,10)$ and $r(v^{\prime})=(3,10,m)$ with $m>10$ (Figure 14), and
-$A_{v,v}^{10}=4227>4225=A_{v^{\prime},v^{\prime}}^{10}.$
-This shows that $r(v)\equiv(3,10,10)$, and the graph is isomorphic to the $(3,10,10)$-solid.
-Figure 14. 5-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(3,10,10)$ and $r(v^{\prime})=(3,10,m)$ with $m>10$.
-Case (6)
-($d=3,r_{1}=4,k=1$) Each face with $m\geq 5$ should be surrounded by exact $m/2$ quadrangular faces and thus $m$ is even.
-As $p(v_{1})>\frac{1}{2}$ by (4.1), we must have $r_{2}(v_{1})=6$ and $r_{3}(v_{1})=6,8$ or $10$.
-Sub-case: $r_{3}(v_{1})=6$
-Suppose $r(v)\not\equiv(4,6,6)$.
-Then there exists a quadrangular face with two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,6)$ and $r(v^{\prime})=(4,6,m)$ with $m>6$ (Figure 15).
-By counting closed walks we obtain
-$A_{v,v}^{6}=115>113=A_{v^{\prime},v^{\prime}}^{6}.$
-Thus $r(v)\equiv(4,6,6)$, and the graph is isomorphic to the $(4,6,6)$-solid.
-Figure 15. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,6)$ and $r(v^{\prime})=(4,6,m)$ with $m>6$.
-Sub-case: $r_{3}(v_{1})=8$
-For this we claim that for every vertex $v$ with $r(v)=(4,6,8)$, its adjacent vertex $v^{\prime}$ outside the hexagonal face $f_{2}(v)$ also has $r(v^{\prime})=(4,6,8)$.
-Indeed, if this is violated by some adjacent pair $v,v^{\prime}$ (Figure 16), then
-$A_{v,v}^{6}=113>111=A_{v^{\prime},v^{\prime}}^{6}.$
-This verifies the claim.
-Figure 16. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,8)$ and $r_{2}(v^{\prime})>6$.
-Now suppose $r(v)\not\equiv(4,6,8)$.
-Then by the claim there exist two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,8)$ and $r(v^{\prime})=(4,6,m)$ with $m>8$ (Figure 17).
-A counting of closed walks with length $8$ gives
-$A_{v,v}^{8}=811>809=A_{v^{\prime},v^{\prime}}^{8}.$
-Hence $r(v)\equiv(4,6,8)$, and the graph is isomorphic to the $(4,6,8)$-solid.
-Figure 17. 4-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,8)$ and $r(v^{\prime})=(4,6,m)$ with $m>8$.
-Sub-case: $r_{3}(v_{1})=10$
-Note that any two adjacent vertices $v,v^{\prime}$ with $r(v)=(4,6,10)$ and $r_{2}(v^{\prime})>6$ (Figure 18) satisfy
-$A_{v,v}^{6}=113>111=A_{v^{\prime},v^{\prime}}^{6}.$
-This shows that any vertex that is adjacent to some vertex with label $(4,6,10)$ should lie on a hexagonal face.
-Figure 18. 3-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,10)$ and $r_{2}(v^{\prime})>6$.
-Moreover, if $r(v)\not\equiv(4,6,10)$, then there exist two adjacent vertices $v,v^{\prime}$ such that $r(v)=(4,6,10)$ and $r(v^{\prime})=(4,6,m)$ with $m>10$ (Figure 19), which give
-$A_{v,v}^{10}=6065>6063=A_{v^{\prime},v^{\prime}}^{10}.$
-Hence $r(v)\equiv(4,6,10)$, and the graph is isomorphic to the $(4,6,10)$-solid.
-Figure 19. 5-neighborhood of $\{v,v^{\prime}\}$, where $r(v)=(4,6,10)$ and $r(v^{\prime})=(4,6,m)$ with $m>10$.
-Case (12)
-($d=4,r_{1}=3,k=2$) We start with a lemma.
-Lemma 4.3.
-Let $G$ be $d$-regular with $d\geq 3$. Then,
-$A_{v_{i},v_{i}}^{4}-(2d^{2}-d)=2(\#\text{ of $4$-cycles through $v_{i}$}).$
-Proof.
-Every closed walk of length $4$ is either contractible or traces a $4$-cycle in one of two directions, and there are exactly $d(d-1)+d^{2}$ contractible closed walks of length $4$. ∎
-We understand each vertex is adjacent to two triangles and two other faces, say of degree $4\leq m\leq m^{\prime}$ respectively, both larger than $3$.
-We subdivide the argument into four sub-cases where there are zero, one, two, or more than two $4$-cycles through each vertex.
-Sub-case: Zero 4-cycles
-There are no squares, and no two triangles share an edge.
+($d=4,r_{1}=3,k=2$) We first justify Lemma 4.5 for this case.
+Proof of Lemma 4.5 for case (12).
+Suppose a chordless 5-cycle $C=\overline{a_{1}a_{2}a_{3}a_{4}a_{5}a_{1}}$ is not a face.
+Note that the number of inner incident edges of $C$ must be even.
+By the 3-connectedness, the possible choices for $d_{*}(C)$ are $(0,0,1,1,2)$, $(0,1,0,1,2)$, $(0,1,1,1,1)$, $(0,1,1,2,2)$, $(0,1,2,1,2)$ or $(1,1,1,1,2)$.
+In view of the symmetry between the inside and outside of $C$, it suffices to rule out the first three situations, as illustrated in Figure 9.
+Figure 9. Three impossible 5-cycles with $d_{*}(C)=(0,0,1,1,2)$, $(0,1,0,1,2)$ and $(0,1,1,1,1)$ respectively.
+Sub-case: $d_{*}(C)=(0,0,1,1,2)$.
+Note that one of the two incident triangles of $a_{1}$ is bounded by $\overline{a_{1}a_{2}}$, and one of the two incident triangles of $a_{5}$ is bounded by $\overline{a_{4}a_{5}}$.
+To have three distinct inner neighbors of $C$ (by the 3-connectedness), $a_{3}$ has no incident triangles inside $C$, and thus $a_{2},a_{3},a_{4}$ have a common outer neighbor.
+This leads to $\mathcal{C}_{4}(a_{4})\geq 2>\mathcal{C}_{4}(a_{5})$, a contradiction.
+Sub-case: $d_{*}(C)=(0,1,0,1,2)$.
+As in the first sub-case, a triangle outside $C$ is bounded by $\overline{a_{1}a_{2}}$ and one inside $C$ is bounded by $\overline{a_{4}a_{5}}$.
+It follows similarly that $a_{2}$ has no inner incident triangles, and this forces $a_{1},a_{2},a_{3}$ to have a common neighbor outside $C$ which has three incident triangles.
+Sub-case: $d_{*}(C)=(0,1,1,1,1)$.
+By the 3-connectedness, there is at most one inner incident triangle of $C$.
+We also note that, in $\{a_{2},a_{3},a_{4},a_{5}\}$ the vertices with two incident triangles outside $C$ cannot be adjacent, otherwise four of the vertices on $C$ have a common neighbor that is incident to three triangles.
+Therefore, the only possible choice of the unique inner incident triangle of $C$ is the one bounded by $\overline{a_{3}a_{4}}$.
+However, in this situation $C$ has only two outer neighbors, which violates the 3-connectedness.
+∎
+Now we are ready to finish off Case (12).
+In this case, each vertex is incident to two triangles and two other faces of length $\geq 4$.
+In view of Lemma 4.1, we subdivide the argument into four sub-cases where $\mathcal{C}_{4}(v)\equiv\mathcal{C}_{4}=0,1,2$ or $\geq 3$.
+Sub-case: $\mathcal{C}_{4}=0$
+There are no faces of length $4$, and no two triangles share an edge.
+It follows that for each $v$, $r(v)$ is of the form $(3,m,3,m^{\prime})$ for some $5\leq m\leq m^{\prime}$.
 Such graphs have plenty of structure.
-Firstly, each face of $\geq 5$ sides is adjacent only to triangular faces.
-Furthermore, each such face is, through each vertex, opposed to another face of $\geq 5$ sides.
-We first establish that, in this situation,
-$A_{v_{i},v_{i}}^{5}-60=2\times(\#\text{ of pentagons incident to $v_{i}$})$
-Our job is now to eliminate the cases where
-$r(v)=(3,5,3,6)\text{ or }(3,5,3,7)$
-for each $v$, and we will be left with $r(v)\equiv(3,5,3,5)$.
-In fact, if we have either case above for one vertex, we have it for all of them.
-We can construct a $3$-neighborhood of the vertex $v$ (Figure 20) with either a hexagon or a face of $\geq 7$ sides and obtain
-$A_{v,v}^{6}=270+\begin{cases}2&r(v)=(3,5,3,6)\\ 0&r(v)=(3,5,3,7).\end{cases}$
-Figure 20. Two possible 3-neighborhood of $v$.
-If we attempt to construct either graph with either $r\equiv(3,5,3,6)$ or $r\equiv(3,5,3,7)$, we quickly run into absurdities. Hence, we are left with the $(3,5,3,5)$ solid.
-Sub-case: one $4$-cycle
+Firstly, each face of length $\geq 5$ is adjacent only to triangular faces.
+Furthermore, each such face is, through each vertex, opposed to another face of length $\geq 5$.
+Using Lemma 4.1 and (4.1), we have $\mathcal{C}_{5}(v)=\mathcal{C}_{5}(v_{1})\geq 1$ for any vertex $v$.
+Also we note that every 5-cycle is chordless, and so is a face by Lemma 4.5.
+If we attempt to construct a triangle on which each vertex is incident to exactly one pentagon face, we quickly run into absurdities.
+Hence, we are left with $r(v)\equiv(3,5,3,5)$, and the graph is isomorphic to the $(3,5,3,5)$-solid.
+Sub-case: $\mathcal{C}_{4}=1$
 First, we rule out the possibility that such a graph has a square face.
@@
 Figure 10. Our “bee” graph.
-Note, one or both of the antennae may attach to a wing to form a pentagon. Running some software, we find that, for any vertex $v$ of the square,
-$A_{v,v}^{5}=\begin{cases}68&\text{ if no pentagon is incident to $v$}\\ 70&\text{ if the remaining face incident to $v$ is a pentagon}.\end{cases}$
-On the other hand, the vertex $v^{\prime}$ at the “neck” of the bee satisfies
-$A_{v^{\prime},v^{\prime}}^{5}=62+2\times(\#\text{ of pentagons incident to $v^{\prime}$})\leq 66,$
-and so we have $A_{v^{\prime},v^{\prime}}^{5}<A_{v,v}^{5}$, a contradiction.
-We have shown that the only $4$-cycles we find are the boundary of the union of two triangular faces.
-Since there are no squares, there must be at least one pentagon.
-The graph self-destructs upon any attempt to draw a $1$-neighborhood of the pentagon.
-Sub-case: two $4$-cycles
+Note, one or both of the antennae may attach to a wing to form a pentagon face.
+Applying Lemma 4.5, we find that, $\mathcal{C}_{5}(v)\geq 4$ for any vertex $v$ on the square face, while $\mathcal{C}_{5}(v^{\prime})\leq 3$ for the vertex $v^{\prime}$ at the “neck” of the bee.
+This is contrary to Lemma 4.1.
+Combining this and Lemma 4.4, we have shown that the only $4$-cycles we find are the boundary of the union of two triangular faces.
+Since there are no square faces, there must be at least one pentagon face.
+The graph self-destructs upon any attempt to draw the $1$-neighborhood of a pentagon face.
+Sub-case: $\mathcal{C}_{4}=2$
 There are a few situations we must rule out.
@@
 The first is incident to only one $4$-cycle, so is ruled out.
-We try to draw the graph starting at this vertex and obtain the one in Figure 22, which forces a vertex to have at least three $4$-cycles.
+We try to draw the graph starting at this vertex and obtain the one in Figure 11, which forces a vertex to lie on at least three 4-cycles.
 Figure 11. The left-center vertex is forced to be incident to three $4$-cycles.
@@
 Hence, every vertex of our graph must have label $(3,4,3,4)$, and we obtain the $(3,4,3,4)$-solid.
-Sub-case: More than two $4$-cycles
+Sub-case: $\mathcal{C}_{4}\geq 3$
 The only way a vertex can be incident to four $4$-cycles is if its $2$-neighborhood is that in Figure 12. However, this vertex’s neighbors are only ever incident to three $4$-cycles.
 Figure 12. $2$-neighborhood of a vertex incident to four $4$-cycles.
-Hence, all that remains is to exclude the case where the vertices are incident to exactly three $4$-cycles.
-A similar argument as the one above excludes the case where $v$ is incident to a face of $5$ or more sides.
+Hence, all that remains is to exclude the case $\mathcal{C}_{4}=3$.
+A similar argument as the one above excludes the case where $v$ is incident to a face of length $\geq 5$.
 Hence, for each vertex $v$, we have either $r(v)=(3,4,3,4)$ or $r(v)=(3,3,4,4)$.
@@
 Figure 13. The graph with exactly three $4$-cycles through each vertex.
-This graph, however, is not walk-regular. The three innermost and outermost vertices all have $74$ returning walks of length $5$, while the other vertices have $76$.
+This graph, however, is not walk-regular. The three innermost and outermost vertices all have seven incident $5$-cycles, while the other vertices have eight.
 Case (13)
-($d=4,r_{1}=3,k=1$) As in the proof of the previous case, we subdivide the argument according to the number of 4-cycles through each vertex.
-Since there is exactly 1 triangle face adjacent to each vertex, the number of 4-cycles is determined by the number of adjacent squares.
-Therefore, each vertex must be adjacent to the same number of squares.
-By (4.1), we must have $p(v_{1})>1$, and thus there must be at least two squares adjacent to each vertex.
-Sub-case: three squares at each vertex
+($d=4,r_{1}=3,k=1$) As in the proof of the previous case, we subdivide the argument according to $\mathcal{C}_{4}$, the number of 4-cycles through each vertex.
+Since there is exactly 1 triangular face incident to each vertex, $\mathcal{C}_{4}$ is exactly the number of square faces around it.
+By (4.1), we must have $p(v_{1})>1$, and thus $\mathcal{C}_{4}\geq 2$.
+Sub-case: $\mathcal{C}_{4}=3$
 This case is straightforward since we already know that $r(v)\equiv(3,4,4,4)$.
-We remark that, in this case, when one wants to reconstruct the $(3,4,4,4)$-solid, there are two possible graphs, the other one being the twisted $(3,4,4,4)$-solid graph.
-However, there are vertices on the twisted $(3,4,4,4)$ solid that have different numbers of closed walks of length 7, and thus, this twisted case should be ruled out.
+We remark that, however, when one wants to reconstruct the $(3,4,4,4)$-solid, there are two possible graphs, the other one being the twisted $(3,4,4,4)$-solid graph.
+Nonetheless, there are vertices on the twisted $(3,4,4,4)$-solid that have different numbers of closed walks of length 7, and thus, this twisted case should be ruled out.
 See Figure 14.
 Figure 14. (3,4,4,4)-solid and twisted (3,4,4,4)-solid.
-Sub-case: two squares at each vertex
-In this case, $r(v_{1})$ is either $(3,4,4,5)$ or $(3,4,5,4)$.
+Sub-case: $\mathcal{C}_{4}=2$
+In this sub-case, $r(v_{1})$ is either $(3,4,4,5)$ or $(3,4,5,4)$.
 We first consider the case $r(v_{1})=(3,4,4,5)$.
@@
 Here, vertex $v_{1}$ is what we started with.
-We claim that the unknown polygon adjacent to vertex $v_{3}$ must be a pentagon.
-Then a similar argument will show that all outer polygons must be pentagons.
-To do so, we count the closed walks of length 5 for $v_{1}$ and for $v_{3}$.
-Non-trivial ones are either a pentagon or a triangle attached to a square, which we shall call a “house”.
-$v_{1}$ and $v_{3}$ always share the same number of “houses”, and we are done with the claim.
-Next, we claim that we will get a contradiction if the center polygon is not a pentagon.
-In this case, for $v_{1}$ and $v_{2}$ to have the same number of closed walks of length 5, we cannot create more “houses” in the outer layer for vertex $v_{1}$.
-This forces us to get 3 squares in a row.
-After drawing all the squares we arrive at Figure 27, which is impossible since there are vertices adjacent to more than 2 triangles.
-Figure 27. An impossible graph when the center is not a pentagon.
-Therefore, we conclude that we must have a pentagon at the center, and this in turn implies that we must add an additional “house” to vertex $v_{1}$, so we arrive at Figure 28.
+We claim that the unknown face incident to vertex $v_{3}$ must be of length 5.
+Then a similar argument will show that all unknown faces within the 2-neighborhood of the center face must be of length 5.
+To do so, we count the 5-cycles incident to $v_{1}$ and to $v_{3}$.
+Note that $v_{1}$ and $v_{3}$ always share the same number of “houses” (i.e., the 5-cycle with a chord), and $v_{1}$ is already incident to a pentagon face.
+To satisfy $\mathcal{C}_{5}(v_{1})=\mathcal{C}_{5}(v_{3})$, we must have a chordless 5-cycle $C$ through $v_{3}$, and there are three possible situations: (i) $C=\overline{u_{1}u_{2}v_{3}u_{3}u_{4}u_{1}}$, (ii) $C=\overline{u_{2}v_{3}u_{3}u_{4}u_{5}u_{2}}$, and (iii) $u_{2},u_{4}$ have a common neighbor $w$ such that $C=\overline{wu_{2}v_{3}u_{3}u_{4}w}$.
+Among these situations, (i) is simply ruled out as $\{u_{2},u_{4}\}$ would be a cut-set of the graph.
+For (ii), the 3-connectedness yields that each of $u_{2}$ and $u_{5}$ has an incident edge inside $C$, and these two edges must have a common endpoint $x$ to bound the incident triangle of $u_{2}$.
+However, there is still a cut-set $\{x,u_{4}\}$ of the graph.
+Now we are left with (iii).
+Using a similar argument, it is impossible to have $\geq 3$ inner neighbors of $C=\overline{wu_{2}v_{3}u_{3}u_{4}w}$, which is a contradiction.
+Hence, this $C$ must be a face with length $5$, and thus the claim is justified.
+Next, we claim that we will get a contradiction if the center face is of length $\geq 6$.
+In this case, to have $\mathcal{C}_{5}(v_{1})=\mathcal{C}_{5}(v_{2})$, we cannot create more “houses” in the outer layer for vertex $v_{1}$.
+This forces us to get three square faces in a row.
+After drawing all the square faces we arrive at Figure 16, which is impossible since there are vertices incident to more than 2 triangles.
+Figure 16. An impossible graph when the center face is of length $\geq 6$.
+Therefore, we conclude that we must have a pentagon face at the center, and this in turn implies that we must add an additional “house” to vertex $v_{1}$, so we arrive at Figure 17.
 Figure 17. 3-neighborhood of $v_{1}$ with $r(v_{1})=(3,4,4,5).$
 We remark that this is exactly part of the twisted $(3,4,5,4)$-solid.
-To get a contradiction, we count the number of closed walks of length 6 passing through $v_{1}$ and through $v_{2}$.
-The non-trivial shapes that contribute to the number of closed walks of length 6 are graphed in Figure 29.
-Figure 29. Possible non-trivial shapes contributing to the number of closed walks of length 6.
-One can see that $v_{1}$ and $v_{2}$ share the same count for all shapes except for the double-square, for which $v_{1}$ has 1 but $v_{2}$ has 2, a contradiction.
+Counting the number of 6-cycles through $v_{1}$ and through $v_{2}$, we have $\mathcal{C}_{6}(v_{1})=4\neq 5=\mathcal{C}_{6}(v_{2})$, which is contrary to Lemma 4.1.
 Therefore, there is no vertex $v$ satisfying $r(v)=(3,4,4,5)$ in this sub-case, and $v_{1}$ satisfies $r(v_{1})=(3,4,5,4)$.
-Again by counting closed walks of length 5, one can easily show that all vertices adjacent to $v_{1}$ must be adjacent to a pentagon as well, and as a consequence, $r(v)\equiv(3,4,5,4)$.
-The proof is complete.
+By a similar counting of the 5-cycles, we exclude $r(v)=(3,4,4,m)$ with $m>5$, as such vertex $v$ gives $\mathcal{C}_{5}(v)\leq 4<\mathcal{C}_{5}(v_{1})$.
+Now the label of each $v$ is of the form $(3,4,m,4)$ with $m\geq 5$.
+Again, counting the 5-cycles in the 2-neighborhood of $v$, we have
+$\mathcal{C}_{5}(v)=\begin{cases}\,6\quad\text{if }r(v)=(3,4,5,4),\\ \,5\quad\text{if }r(v)=(3,4,m,4)\text{ with }m>5.\end{cases}$
+As a consequence, $r(v)\equiv(3,4,5,4)$, and we obtain the $(3,4,5,4)$-solid.
+We remark that the above discussion also completes the proof of Lemma 4.5 for $d=4$.
+4.4. The 5-regular cases
+Finally we assume that $G$ is 5-regular. The following discussion completes the proof of Lemma 4.3.
+Proof of Lemma 4.3 for $d=5$.
+Suppose that a triangle $C=\overline{a_{1}a_{2}a_{3}a_{1}}$ is not a face.
+The 3-connectedness implies that $d_{*}(a_{i})\in\{1,2\}$ for all $i$.
+Again, we further assume that such $C$ is minimal among all the triangles that are not faces.
+Then each $v$ inside $C$ cannot be adjacent to all vertices $a_{1},a_{2},a_{3}$ of $C$, and around such $v$ there is at most one face with length $\geq 4$ since $\mathcal{C}_{3}(v)=\mathcal{C}_{3}(v_{1})\geq 4$.
+Moreover, it follows that each edge inside $C$ lies on at least one triangle.
+In view of these facts, we can simply rule out the case $d_{*}(C)=(1,1,1)$, and eliminate the other three cases by drawing the adjacent vertices of $C$:
+Case: $d_{*}(C)=(1,1,2)$.
+Note that the two adjacent vertices $b_{1},b_{2}$ of $a_{1},a_{2}$ inside $C$ cannot be identical, otherwise around this common neighbor there are at least two faces with length $\geq 4$.
+Since each of the edges $\overline{a_{1}b_{1}},\overline{a_{2}b_{2}}$ belongs to some triangle, both $b_{1}$ and $b_{2}$ must be adjacent to $a_{3}$ (Figure 18).
+Now $\{b_{1},b_{2}\}$ is a cut-set of the graph, and this violates the 3-connectedness.
+Figure 18. An impossible triangle with $d_{*}(C)=(1,1,2)$.
+Case: $d_{*}(C)=(1,2,2)$.
+Let $b_{1}$ be the neighbor of $a_{1}$ inside $C$.
+As $\mathcal{C}_{3}(a_{1})\geq 4$ and $C$ is minimal, we can see that the edge $\overline{a_{1}b_{1}}$ belongs to exactly one triangle, and there are at least two incident triangles of $a_{1}$ outside $C$.
+Therefore, one of $a_{2},a_{3}$ (say $a_{2}$) has a common neighbor with $a_{1}$ outside $C$.
+By the 3-connectedness there are at least three distinct neighbors of $\{a_{1},a_{2},a_{3}\}$ outside $C$, and thus $a_{3}$ has no common neighbors with $a_{1}$ and $a_{2}$ outside $C$.
+Now we have $\mathcal{C}_{3}=4$, which forces that $a_{2},a_{3}$ have exactly two and three incident triangles inside $C$ respectively.
+We can draw four incident triangles $\overline{a_{1}b_{1}a_{3}a_{1}},\overline{a_{2}b_{3}b_{2}a_{2}},\overline{a_{2}a_{3}b_{3}a_{2}}$ and $\overline{a_{3}b_{1}b_{3}a_{3}}$ as in Figure 19.
+As $\mathcal{C}_{3}(b_{1})=\mathcal{C}_{3}(b_{2})=4$, the undetermined neighbor of $b_{3}$ must be adjacent to $b_{1}$ and $b_{2}$.
+This leads to $\mathcal{C}_{3}(b_{3})=5$, a contradiction.
+Figure 19. An impossible triangle with $d_{*}(C)=(1,2,2)$.
+Case: $d_{*}(C)=(2,2,2)$.
+Again by the 3-connectedness, to have three distinct neighbors of $\{a_{1},a_{2},a_{3}\}$ outside $C$, no incident triangles outside $C$ are allowed.
+It follows from $\mathcal{C}_{3}\geq 4$ that each $a_{i}$ has exactly three incident triangles inside $C$ as in Figure 20.
+However, the center triangle $\overline{b_{1}b_{2}b_{3}b_{1}}$ cannot be a face, and this violates the minimality of $C$.
+The proof of Lemma 4.3 is complete.
+Figure 20. An impossible triangle with $d_{*}(C)=(2,2,2)$.
+As a result, in case (14) we have $r(v)\equiv(3,3,3,3,3)$, and the graph is isomorphic to icosahedron.
+To solve the last case ($d=5,r_{1}=3,k=4$), we also need the fact that every chordless 4-cycle must be a face there.
+Proof of Lemma 4.4 for $d=5$.
+Suppose that a chordless $4$-cycle $C=\overline{a_{1}a_{2}a_{3}a_{4}a_{1}}$ is not a face, and such $C$ is minimal.
+This means that every chordless $4$-cycle inside $C$ must be a face, and thus, every vertex inside $C$ is adjacent to at most $2$ vertices on $C$.
+Note that $\mathcal{C}_{3}=\mathcal{C}_{3}(v_{1})=4$ or $5$.
+If $\mathcal{C}_{3}=5$, then each $a_{i}$ has exactly two neighbors inside $C$, and $\{a_{1},a_{2},a_{3},a_{4}\}$ must have a common neighbor outside $C$, which is a cut-vertex of the graph.
+Therefore, we are left with $\mathcal{C}_{3}=4$.
+For the case $d_{*}(a_{1})=0$, it is impossible to have $\geq 3$ distinct neighbors of $\{a_{2},a_{3},a_{4}\}$ inside $C$, which violates the 3-connectedness.
+Similarly, we rule out the case $d_{*}(a_{i})=3$.
+Since the inner degree sequence is defined lexicographically, if $d_{*}(a_{1})=2$, then we must have $d_{*}(C)=(2,2,2,2)$.
+Since there are at least three distinct neighbors outside $C$, the four incident edges outside $C$ provide at most one incident triangle, which is not enough to make $\mathcal{C}_{3}(a_{i})=4$ for each $i=1,2,3,4$.
+Now the only possibility is $d_{*}(a_{1})=1$.
+Let $b_{1}$ be the neighbor of $a_{1}$ inside $C$.
+To have $\mathcal{C}_{3}(a_{1})=4$, $a_{1}$ has exactly three incident triangles outside $C$, and either $a_{2}$ or $a_{4}$ should be the second neighbor of $b_{1}$ on $C$.
+As the only face around $b_{1}$ with length $>3$ is bounded by $\overline{a_{1}b_{1}}$, the second neighbor of $b_{1}$ on $C$ should have two neighbors inside $C$.
+Since $d_{*}(a_{2})\leq d_{*}(a_{4})$, we may assume that this neighbor of $b_{1}$ is $a_{4}$, and $d_{*}(a_{4})=2$.
+If $d_{*}(a_{2})=1$, then following the same argument as above, the neighbor $b_{2}$ of $a_{2}$ inside $C$ is adjacent to $a_{3}$, and $d_{*}(a_{3})=2$.
+To have $\mathcal{C}_{3}(a_{2})=\mathcal{C}_{3}(a_{3})=\mathcal{C}_{3}(a_{4})=4$, there are three incident triangles of $a_{2}$ outside $C$, three incident triangles of $a_{3}$ as well as three of $a_{4}$ inside $C$, and thus $a_{3},a_{4}$ have a common neighbor $b_{3}$ inside $C$ (Figure 21).
+Now either of the two undetermined faces around $b_{3}$ is a triangle.
+However, no matter which one is not a triangle, it becomes the second face around $b_{1}$ or $b_{2}$ with length $>3$, which is contrary to $\mathcal{C}_{3}(b_{1})=\mathcal{C}_{3}(b_{2})=4$.
+Figure 21. An impossible 4-cycle with $d_{*}(C)=(1,1,2,2)$.
+The last case we will rule out is $d_{*}(a_{2})=2$.
+As $\mathcal{C}_{3}(a_{2})=4$, $a_{2}$ has two incident triangles outside $C$, so that the common neighbor of $a_{1}$ and $a_{2}$ is adjacent to $a_{3}$.
+To have three distinct neighbors of $\{a_{1},a_{2},a_{3},a_{4}\}$ outside $C$, $d_{*}(a_{3})$ should be $1$, and hence no triangles are bounded by $\overline{a_{3}a_{4}}$, which violates $\mathcal{C}_{3}(a_{3})=\mathcal{C}_{3}(a_{4})=4$.
+The proof of Lemma 4.4 is complete.
+∎
+Case (15)
+($d=5,r_{1}=3,k=4$) As $p(v_{1})>\frac{3}{2}$ by (4.1), $r_{5}(v_{1})=4$ or $5$.
+Since around each vertex four triangular faces are determined by Lemma 4.3, we are able to draw the 2-neighborhood of a face with length $\geq 4$, as in Figure 22.
+It is straightforward to count that
+$\mathcal{C}_{4}(v)=\begin{cases}\,7\quad\text{if }r(v)=(3,3,3,3,4),\\ \,6\quad\text{if }r(v)=(3,3,3,3,m)\text{ with }m>4.\end{cases}$
+In view of Lemma 4.1, $r(v)\equiv(3,3,3,3,4)$ and the graph is isomorphic to the $(3,3,3,3,4)$-solid provided that $r_{5}(v_{1})=4$.
+Figure 22. The $2$-neighborhood of faces with length $4$ and $>4$ respectively.
+When $r_{5}(v_{1})=5$, each vertex $v$ lies on a face with length $\geq 5$.
+Note that in the 2-neighborhood of such face, the vertices in the 2-neighborhood of $v$ with undetermined neighbors are $u_{1}$, $u_{2}$ and $u_{3}$.
+As there are no faces with length $4$, $u_{2}$ is not adjacent to $u_{3}$.
+Moreover, $u_{1}$ is not adjacent to $u_{3}$ either, otherwise they form a 4-cycle that is not a face, which is contrary to Lemma 4.4.
+Therefore, each 5-cycle through $v$ can be found in the Figure 23, which is either a union of three adjacent triangles or a face with length $5$ (we remark that the proof of Lemma 4.5 is complete here).
+A counting of these 5-cycles gives
+$\mathcal{C}_{5}(v)=\begin{cases}\,11\quad\text{if }r(v)=(3,3,3,3,5),\\ \,10\quad\text{if }r(v)=(3,3,3,3,m)\text{ with }m>5.\end{cases}$
+Again by Lemma 4.1, we finally have $r(v)\equiv(3,3,3,3,5)$, and the graph is isomorphic to the $(3,3,3,3,5)$-solid.
+Figure 23. The $2$-neighborhood of faces with length $5$ and $>5$ respectively when $r_{5}(v_{1})=5$.